### v1

# Contrasting exchange-field and spin-transfer torque driving mechanisms in all-electric electron spin resonance

Abstract
Understanding the coherent properties of electron spins driven by electric fields is crucial for their potential application in quantum-coherent nanoscience. In this work, we address two distinct driving mechanisms in electric-field driven electron-spin resonance as implemented in scanning tunneling spectroscopy. We study the origin of the driving field using a single orbital Anderson impurity, connected to polarized leads and biased by a voltage modulated on resonance with a spin transition. By mapping the quantum master equation into a system of equations for the impurity spin, we identify two distinct driving mechanisms. Below the charging thresholds of the impurity, electron spin resonance is dominated by a magnetically exchange-driven mechanism or field-like torque. Conversely, above the charging threshold spin-transfer torque caused by the spin-polarized current through the impurity drives the spin transition. Only the first mechanism enables coherent quantum spin control, while the second one leads to fast decoherence and spin accumulation towards a non-equilibrium steady-state. The electron spin resonance signals and spin dynamics vary significantly depending on which driving mechanism dominates, highlighting the potential for optimizing quantum-coherent control in electrically driven quantum systems.
I Introduction
Coherent manipulation of individual electron spins localized in single atoms or molecules on a surface using electron spin resonance (ESR) implemented in a scanning tunneling microscope (STM) is a novel platform for the study of many-body physics with possible applications for quantum simulations using artificially constructed interacting spin systems on surfaces  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Such a system can be modeled as a magnetic impurity connected to two leads, which provide a harmonic alternating current (AC) drive and a static external magnetic field that Zeeman-splits the up and down spin states. When one of the leads is ferromagnetic, a polarized current is detected in the direct current (DC) signal. In this situation, adjusting the driving frequency of the AC voltage to correspond with the Zeeman splitting leads to resonance behavior in the electric current.
The mechanism mediating the coupling between the AC electric field and the spin remains a topic of debate [26, 27, 28, 29, 30, 31]. Recent studies have proposed that the driving mechanism originates from the magnetic field component of the electric field, which is sufficiently amplified in the electromagnetic cavity formed between the tip and the magnetic impurity [30]. This driving is necessary because a time-dependent Kondo scattering Hamiltonian lacks the charge fluctuations required to provide any driving to explain the experimentally observed Rabi rates, as we mention in [27]. The presence of an AC magnetic field as a product of the AC electric field contrasts with recent experimental findings [21, 22, 20], which suggest that a magnetic atom must be placed near the spin impurity to enable driving outside the tunnel junction but still under the influence of the electric field generated by the tip body. A long-range electromagnetic driving field appears inconsistent with the requirement for a localized magnetic field source.
More recently, a Kondo Hamiltonian approach was used to derive an effective magnetic field proportional to the tip’s spin polarization and the AC voltage, which can drive the magnetic impurity [31]. This model accounts for current flow via multi-electron (cotunneling) processes in a region where sequential tunneling is suppressed by Coulomb blockade [32, 33, 27]. While this proposed driving mechanism aligns with the approach used in this work, it does not fully capture all features of the experimentally observed homodyne detection. Homodyne detection arises from the misalignment between the STM tip’s spin and the magnetic impurity and the AC driving, imposing a change on the tunneling magnetoresistance, as observed in ESR measurements. In particular, Homodyne detection predicts the strongest ESR signal when the tip spin polarization and the magnetic impurity are mostly parallel [34, 21], whereas [31] differs from previous reports by showing the maximum ESR signal at $90^{\circ}$.
Some of us have proposed an all-electric model for ESR in the STM, which can efficiently drive the spin of an Anderson impurity model based on the principle of barrier modulation [35, 36]. We were able to solve the resulting quantum master equation either in the long-time limit or explore real-time spin dynamics by solving the time propagation of the Hamiltonian [37]. At present, the model is limited to the weak coupling regime, imposing sequential tunneling processes, which implies no cotunneling effects, although the model is extended to the Coulomb Blockade regime via the tail of the resonance [36]. It is also important to note that much of the existing literature on cotunneling in transport overlooks the exchange field contribution from the sequential regime. As pointed out by Braun et al. in Refs. [38, 39], this approach is inaccurate. A comprehensive theory that fully integrates the main elements of both the sequential and cotunneling regimes for a periodically driven system is still lacking, but recent works are aiming towards this goal [40, 41, 42]. Nevertheless, our extended sequential model qualitatively describes many of the experimentally observed features of ESR of single spin impurities on surfaces.
However, a fundamental question was left unanswered: how does the barrier modulation couple to the spin, and can this mechanism be linked to any physical measurements? In this work, we aim to explore this question by establishing connections with related models for quantum dots connected to magnetic leads [43, 44, 45, 38, 42] where an exchange field mechanism between the tip and sample is demonstrated and applied to experimental studies [46, 47]. This comparison is justified, as atomic or molecular spin impurities can be seen as miniaturized versions of lithographically defined larger quantum dots. We will show that the ESR mechanism is highly dependent on the applied DC bias relative to the ionization and charging energy thresholds of the impurity i.e., it depends if we are above the energy thresholds or inside the Coulomb Blockade regime. Specifically, we will contrast two driving mechanisms:
  • Field-like Torque which dominates ESR for DC biases in the Coulomb blockade regime. It is mediated by the exchange magnetic field induced by the spin-polarized tip, leading to a conventional ESR mechanism and enabling coherent spin manipulation via electric pulses. We refer to this driving mechanism as FLT.
  • Spin-transfer Torque which drives ESR for DC biases outside the Coulomb blockade regime. It is dominated by spin accumulation on the impurity, driven by the current which torques the impurity spin and leads to strong spin polarization of the spin impurity. We refer to this driving mechanism as STT.
We emphasize that the driving mechanism is a distinct concept and exists separately from the transport regime, which is determined by the DC bias value relative to the ionization energies. The driving mechanisms (FLT, STT) are general and exist below and above the ionization thresholds. The transport regime dictates how strongly FLT and STT affect the magnetic impurity.
This paper is organized as follows. In Sec. II, we summarize the theoretical approach, drawing the comparison to quantum dot models. In Sec. III, we analyze the key physical findings from this comparison, defining the driving regimes and how they affect the quantum impurity. Finally, we perform numerical simulations to verify the key features of the theory, leading to the conclusions.
II Transport Model
II.1 Model Hamiltonian
The model used here is a single orbital Anderson impurity model [48] (SAIM) of a quantum impurity (QI), $H_{\mathrm{QI}}$, connected to two ferromagnetic or normal electrodes (tip or substrate) that can be treated as fermionic baths by time-dependent hoppings $H_{T}(t)$ as introduced in previous works[35, 36], see Fig. 1a. This model is general enough to describe a wide range of physical systems, where the impurity spin can be localized in point defects, atoms or molecules. The full system is given by $H(t)=H_{\rm L}+H_{\rm R}+H_{\rm QI}+H_{\rm T}(t),$ where the first two terms describes the spin polarized or normal fermionic baths, the second term is the QI and the third term contains the hopping connecting baths and QI. The bath itself can be expressed using the common form (Eq. 1), where each state is described by an electrode label $\alpha=L,R$, spin projection along the $z$ axis $\sigma=\uparrow,\downarrow$, and momentum $k$. For both ferromagnetic electrodes, we consider that their magnetizations point in the $z$ direction. In this case, the effective spin asymmetry, $D_{\alpha\uparrow}\neq D_{\alpha\downarrow}$, in the spin-dependent density of states $D_{\alpha\sigma}$ at the Fermi level can be described by the spin polarization $P_{\alpha}\equiv(D_{\alpha\uparrow}-D_{\alpha\downarrow})/(D_{\alpha\uparrow}+D_{\alpha\downarrow})$, where $D_{\alpha\sigma}=D_{\alpha}\left(1/2+\sigma P_{\alpha}\right)$ with $P_{\alpha}\in[-1,1]$ and the extreme cases representing a fully spin polarized lead pointing down, $-1$, or up, $+1$.
For the sake of simplicity, we focus on the simplest QI in Eq. 2 which contains the Coulomb repulsion $U$, the onsite energy $\varepsilon$ and an external magnetic field $\mathbf{B}=B(\sin\phi\sin\theta,\cos\phi\sin\theta,\cos\theta)$ expressed in spherical coordinates pointing in any direction, noncollinear with the direction of ferromagnetic lead magnetizations. Thus [^1] Heisenberg exchange coupling $J$ between transport spin $\mathbf{\hat{s}}$ and another localized spin or between localized spins can be seamlessly incorporated into the model [35], enabling the study of arbitrary spin clusters and more complex spin structures. Future work will explore this in depth., we have
$H_{\rm\alpha}$ $=$ $\sum_{k\sigma}\varepsilon_{\alpha k}c^{\dagger}_{\alpha k\sigma}c_{\alpha k\sigma},$ (1)
$H_{\rm QI}$ $=$ $\sum_{\sigma}\varepsilon d^{\dagger}_{\sigma}d_{\sigma}+U\hat{n}_{d\uparrow}\hat{n}_{d\downarrow}+g\mu_{B}\mathbf{{B}}\cdot\mathbf{\hat{S}}$ (2)
$H_{T}(t)$ $=$ $\sum_{\alpha k\sigma}\left(t_{\alpha}(t)c^{\dagger}_{\alpha k\sigma}d_{\sigma}+t_{\alpha}^{*}(t)d^{\dagger}_{\sigma}c_{\alpha k\sigma}\right).$ (3)
Each electrode is characterized by a chemical potential $\mu_{\alpha}$ and a temperature $T_{\alpha}$. Our convention for the applied bias is such that $eV_{DC}=\mu_{L}-\mu_{R}$, meaning the DC bias is applied to the left electrode which we will assign as the tip, see Fig. 1a and b. Consequently, electrons move from sample to tip at positive applied bias, with the SOMO energy being associated with $\varepsilon$ and the SUMO energy with $\varepsilon+U$, where $\varepsilon$ and $U$ are the ionization and Coulomb repulsion energies respectively. Changing the sign convention to one where the DC bias is applied to the sample, $eV_{DC}=\mu_{R}-\mu_{L}$, only requires swapping the labels for the $\varepsilon$ and $\varepsilon+U$ energies to keep the correct physical meaning of the orbitals. The right chemical potential is assumed to be nearly aligned with the Fermi energy and grounded, which goes in line with the asymmetry tunneling conditions $t_{L}>t_{R}$ in [49] and follows the experimental evidence of the impurity being more strongly coupled to the substrate than the STM tip. In this model, only the hopping terms include a periodic modulation in the form:
$t_{\alpha}(t)=t_{\alpha}^{0}\left[1+A_{\alpha}\cos(\omega t)\right],$ (4)
with $A_{\alpha}\propto V_{\mathrm{AC}}$. This functional dependence has already been analyzed in previous works [27, 35, 36], so we defer a more detailed examination to a future publication. It is important to highlight that this approach does not explicitly include nor exclude the piezoelectric effect [50]; rather, it accounts for all possible physical contributions on the driving parameter.
Figure 1: a) Schematic of a QI connected a left (L) and right (R) lead with a bias voltage applied to the left lead. We consider the impurity to be more strongly coupled to the R lead. b) Energy scheme of the QI. The hollow arrows represent unoccupied levels of the QI at $\varepsilon+U$ as a product of charging the QI with one more electron to be doubly occupied with a total energy $E_{2}=2\varepsilon+U$. Meanwhile, the filled arrows, which defined the spin of the QI, are occupied single-electron states at $\varepsilon$. Only the left electrode is polarized in this example ($P_{L}>0$) and modulated ($t_{L}=t_{L}(t)$) being $t_{L}<t_{R}$. The applied positive bias ($eV_{\mathrm{DC}}=\mu_{L}-\mu_{R}>\varepsilon+U>0$) implies that the Rabi process $\Omega$ is dominated by STT, as for when ($eV_{\mathrm{DC}}<\varepsilon<0$). Contrary, if $\varepsilon<eV_{\mathrm{DC}}<\varepsilon+U$, the Rabi process is FLT-driven, as indicated on the right side of the figure. The transport through the QI in b)is the following: 50% of electrons up and down will tunnel from the right lead, overcoming the $\varepsilon+U$ energy, forcing the QI to be doubly populated. Since tip accepts more up electrons, the population of the up state in $\varepsilon+U$ will be depleted, leaving it down populated. Since spin-down SUMO state corresponds to a spin-up SOMO state, the QI ends up populated. The full process provides an incoherence resonance in the current when the L tunneling is modulated.c) Convention for the magnetic field direction with respect to the external coordinate system with the azimuthal angle $\phi=0$.
The use of a reduced density matrix formalism allows us to define any observable straightforwardly. To do so, we consider all possible charge configurations: zero, one, and two electrons in the QI (see Fig. 1b). These configurations are denoted as $|p\rangle$, with $p=g,e,0,2$. The first two are the singly occupied states of the QI at $\varepsilon$ energy, corresponding to the ground and excited states, respectively, and their time evolution fully describes the spin state of the system. The states 00 (empty) and $2$ (doubly occupied) are transient states necessary to facilitate transport through the impurity. The first state represents an empty quantum dot, which is aligned with the Fermi energy and therefore chosen to be at zero energy, $E_{0}=0$. Consequently, removing one electron from the impurity requires an energy cost of $E_{1}-E_{0}=\varepsilon$. Meanwhile, the doubly occupied state is formed when two electrons with opposite spins occupy the same level or orbital, with an energy given by $E_{2}=2\varepsilon+U$. To create this state, an electron from the leads must overcome an energy barrier of $E_{2}-E_{1}=\varepsilon+U>0$ to be placed at the singly occupied state. From the energy perspective, the quantum impurity hosts two electrons, one at energy $\varepsilon$ and another at $\varepsilon+U$, with the latter being a effective unoccupied singly state. Thus, the relevant energies for transport are the differences between states with different numbers of electrons, i.e., $\varepsilon$ and $\varepsilon+U$, as shown in Fig. 1b, rather than the absolute energy of the SIAM levels.
Since the impurity Hamiltonian does not depend on time, it is convenient to diagonalize the QI Hamiltonian and introduce the reduced density matrix. Thus, we can write $H_{\rm QI}|l\rangle=\sum_{j}E_{j}\hat{\rho}_{jj}|l\rangle=E_{l}|l\rangle.$ The Hubbard operators $\hat{\rho}_{lj}=|l\rangle\langle j|$, with the Latin characters ($l,j$ and later $u,v,\dots$) referring to eigenstates of the impurity, are closely related to the reduce density matrix [35, 36, 51]. The specific case of the eigenstates $g$ and $e$ to be $\downarrow$ and $\uparrow$, respectively, in Fig. 1b corresponds to a positive out-of-plane magnetic field, i.e., along the positive $z$-axis. As we will later mention, the external magnetic field forces the spin to be aligned parallel to it for an isotropic $g$-factor. This implies that the spin will be quantized in the magnetic field direction rather than $z$, and so the reduced density matrix, by definition, is represented in the coordinate system, associated with the magnetic field direction, where the spin is quantized and the QI Hamiltonian is diagonal. Thus, for example, the expectation value of the spin components needs to be computed with care, see Appendix A, since we cannot simply use the Pauli matrices.
The Hubbard operators will emerge in the hopping Hamiltonian as well as [52],
$H_{T}(t)=\sum_{\alpha k\sigma lj}\left(t_{\alpha}(t)c^{\dagger}_{\alpha k\sigma}\lambda_{lj,\sigma}\hat{\rho}_{lj}+h.c.\right).$
The matrix element $\lambda_{lj,\sigma}=\langle l|d_{\sigma}|j\rangle=\langle l|0\rangle\langle\sigma|j\rangle+\langle l|\bar{\sigma}\rangle\langle 2|j\rangle$ reflects the change of the many-body configurations on the QI, they are computed in Appendix B.
II.2 Quantum master equation
At this stage, we derive a QME for the reduced density matrix by treating the coupling between the QI and the reservoirs up to the lowest order in perturbation theory in $H_{T}$ in the Born-Markov approximation [53, 54], meaning $\hbar^{-1}\Gamma\tau_{B}\ll 1$ where $\Gamma$ is proportional to the rate of change of the reduced density operator due to the coupling to the bath and $\tau_{B}$ is the characteristic decay time of the bath [55]. Following Refs. [56, 57, 58, 59, 60, 61, 51, 35, 36], one arrives at:
$\hbar\dot{\rho}_{lj}(t)$displaystyle-- $i\Delta_{lj}\rho_{lj}(t)=\sum_{vu}\left\{\left[\Gamma_{vl,ju}(t)+\Gamma^{*}_{uj,lv}(t)\right]\rho_{vu}(t)\right.$ (5)
-- $\left.\Gamma_{jv,vu}(t)\rho_{lu}(t)-\Gamma^{*}_{lv,vu}(t)\rho_{uj}(t)\right\},$
where $\Delta_{lj}=E_{l}-E_{j}$. The QME, Eq. 5, is physically meaningful in the limit of weak coupling between QI and electrodes, and, as we mentioned, it is represented in the quantization axis dictated by the spin orientation. In previous works [35], we derived a Floquet version of the QME, which has proven useful in the long-time limit for efficiently computing spectra by transforming the equations to the frequency domain. This formulation incorporates Floquet components of the density matrix, labeled by the Floquet number $n$. The rates in Eq. 5 can be written as the sum of two contributions per electrode $\alpha$:
$\Gamma_{vl,ju}(t)=\sum_{\alpha}\left[\Gamma_{vl,ju,\alpha}^{-}(t)+\Gamma_{vl,ju,\alpha}^{+}(t)\right],$ (6)
which are
$\Gamma_{vl,ju,\alpha}^{+}(t)$ $=$ $\sum_{\sigma}\lambda_{vl,\sigma}\lambda_{uj,\sigma}^{*}\gamma_{\alpha\sigma}(t){\cal I}_{\alpha}^{+}(\Delta_{ju})$ (7)
$\Gamma_{vl,ju,\alpha}^{-}(t)$ $=$ $\sum_{\sigma}\lambda_{lv,\sigma}^{*}\lambda_{ju,\sigma}\gamma_{\alpha\sigma}(t){\cal I}_{\alpha}^{-}(\Delta_{ju}),$ (8)
where the Fermi distribution function $f_{\alpha}(\epsilon)=1/\left(e^{\beta_{\alpha}(\epsilon-\mu_{\alpha})}+1\right)$, being $\beta_{\alpha}=1/k_{B}T_{\alpha}$ and $\mu_{\alpha}$ the chemical potential, is encoded in the integrals ${\cal I}^{\pm}_{\alpha}(\Delta_{ju})$ which are of the form:
[…]
$\times$ $\left(\frac{1}{\epsilon\mp\Delta_{ju}\pm i\hbar/\tau_{c}}+\frac{e^{i\omega t}A_{\alpha}/2}{\epsilon\mp\Delta_{ju}\mp\hbar\omega\pm i\hbar/\tau_{c}}\right.$
$+$ $\left.\frac{e^{-i\omega t}A_{\alpha}/2}{\epsilon\mp\Delta_{ju}\pm\hbar\omega\pm i\hbar/\tau_{c}}\right).$
Notice that if $\hbar\omega\ll|\Delta_{ju}|$, Eq. 9 simplifies to [^2] A feasible assumption since $\hbar\omega$ will be in the 0.1 meV range or below, while $\Delta_{ju}\sim 100$ meV. This is because, in the rates only, $\Delta_{ju}$ always represents an energy difference between charged states with different numbers of electrons; otherwise, the rate is zero due to the definition of $\lambda_{lj,\sigma}$. Conversely, $\Delta_{lj}$ in Eq. 5 can be in the $\hbar\omega$ range.:
[…]
approximation valid for all the simulations in this work. Equation 10 contains $f^{+}_{\alpha}(\epsilon)=f_{\alpha}(\epsilon)$ and $f^{-}_{\alpha}(\epsilon)=1-f_{\alpha}(\epsilon)$ following [38]. Additionally, in Eqs. 7 and 8, we have introduced $\gamma_{\alpha\sigma}(t)=\left(1+A_{\alpha}\cos(\omega t)\right)^{2}\gamma_{\alpha\sigma}^{0}$ where $\gamma_{\alpha\sigma}^{0}=2\pi|t^{0}_{\alpha}|^{2}D_{\alpha}(1/2+\sigma P_{\alpha})=\gamma_{\alpha}^{0}(1/2+\sigma P_{\alpha})$ represents the level broadening due to the time-independent hopping, $t^{0}_{\alpha}$, for spin $\sigma$ in electrode $\alpha$. Further details on the transition rates, the derivation of the quantum master equation (QME), and its solution in the long-time limit can be found in [35].
It is important to mention that $\hbar/\tau_{c}$, up to first order in perturbation theory, should be infinitesimal in Eqs. 9 and 10. However, in this work, $\hbar/\tau_{c}\ll|\Delta_{ju}|\approx|\varepsilon|\ \mathrm{or}\ \varepsilon+U$ but nonzero, allowing us to explore two distinct regimes for the sequential tunneling: the one above the thresholds $|\varepsilon|,\varepsilon+U$ where an full electron can hop in or out from the QI and the one below the thresholds where the QI is in the so called Coulomb blockade (CB) regime and we tunnel through the tail of the resonance [36]. Therefore, the transition between both regimes is controlled by the DC voltages and happens when it equalizes the ionization energy $\varepsilon$ or charging one $\varepsilon+U$, which correspond to the situation of removing and adding one electron on the QI respectively.
We are aware of the limitations a finite $\hbar/\tau_{c}$ introduces, as a proper characterization of the Coulomb blockade (CB) requires the use of cotunneling transport or a Kondo scattering Hamiltonian [39, 27]. However, as long as $\hbar/\tau_{c}\ll|\varepsilon|,\varepsilon+U$, we can overcome this limitation of the model and explore, to some extent, the implications of the FLT regime that governs the ESR signal in the CB regime. In this context, we expect low current at low DC biases, consistent with previous studies [38, 62, 39, 63, 42]. Future extensions of this model will aim to circumvent this issue for a better characterization of the system [^3] Going beyond a linear QME is likely required for this step, making the mathematical treatment of the model more challenging. However, first-order perturbation theory already provides rich results that follow experimental evidence, as we will explore in the next sections.
II.3 Expression for the current
The electric current flowing out of an electrode $\alpha$ is defined following the usual formula [64] as $I_{\alpha}(t)=-e\frac{d\langle N_{\alpha}\rangle}{dt}=-i\frac{e}{\hbar}\langle[H_{T}(t),N_{\alpha}]\rangle$ which leads to:
$I_{\alpha}(t)=\frac{2e}{\hbar}\sum_{lju}\mbox{Re}\left\{\rho_{lu}(t)\left[\Gamma_{lj,ju,\alpha}^{+}(t)-\Gamma_{lj,ju,\alpha}^{-}(t)\right]\right\}.$ (11)
The expression for the current contains both DC and AC terms as a consequence of the hopping modulation. The measurable quantity in the experiment is the electric current signal in the long-time limit, which depends on the resonance condition as well as the relative orientation of the QI spin to the direction of the tip spin polarization. We refer to this current as ESR signal in the following. The coherences (off-diagonal elements of the density matrix) are explicitly included in the expression and are crucial for the correct calculation of the ESR signal, as pointed out in a previous work [36]. The coherences in Eq. 11 also lead to the so called homodyne detection of the electric current, a product of the misalignment of the tip’s spin and the spin of the impurity under it [34, 21, 65]. If the populations of the transient states ($0,2$) are negligible, we arrive at the following expression for the current:
$I_{L}(t)$ $=$ $\frac{e}{\hbar}\gamma_{L}\left(1+A_{L}\cos(\omega t)\right)^{2}$
$\times$ $\left[(1-2P_{L}\langle S_{z}\rangle)\mbox{Re}\left\{{\cal I}_{L}^{+}(\Delta_{2g})\right\}\right.$
-- $\left.(1+2P_{L}\langle S_{z}\rangle)\mbox{Re}\left\{{\cal I}_{L}^{-}(\Delta_{0g})\right\}\right]=-I_{R}(t).$
$I_{L}=-I_{R}$ is due to electric current conservation, more details on the current can be found in Appendix C. In Eq. 12, we have employed the notation $g$ to refer to the ground state for a more general treatment instead of $\downarrow$, accounting for the situation in which the spin is tilted by the magnetic field with respect to the spin polarization on $z$. We do this because the states are always defined relative to the spin’s quantization axis, see Appendix A. A consequence of choosing the $z$-axis as the magnetization direction of the leads, where the spin polarization lies, the electric current becomes a measurement of the expectation value $\langle S_{z}\rangle$ as Eq. 12 shows. In general, for any magnetic orientation of the spin polarized lead, we expect that the current will measure the spin component parallel to it.
Assuming that the magnetic field lies in the $xz$ plane, $\mathbf{B}=B(\sin\theta,0,\cos\theta)$, we have:
$\langle S_{z}\rangle$ $=$ $\frac{\hbar}{2}(\rho_{ee}-\rho_{gg})\cos\theta+\frac{\hbar}{2}(\rho_{ge}+\rho_{eg})\sin\theta.$ (13)
$\theta$, also referred to as homodyne angle, is the angle formed between the direction of the left electrode’s magnetization, fixed along the $z$-axis, and the impurity spin when the $g$-factor is isotropic. We should emphasize that in Eq. 13, the main contribution to the populations comes from the Floquet number $n=0$, while the contribution from the coherences arises from $n=1$ [36]. Thus, the coherences are intrinsically time-dependent in the long-time limit. For the sake of completeness, the in-plane components of the spin expectation value are:
$\langle{S_{x}}\rangle$ $=$ $\frac{\hbar}{2}(\rho_{ee}-\rho_{gg})\sin\theta-\frac{\hbar}{2}(\rho_{ge}+\rho_{eg})\cos\theta,$ (14)
$\langle{S_{y}}\rangle$ $=$ $\frac{i\hbar}{2}(\rho_{ge}-\rho_{eg})=-\hbar\mbox{Im}(\rho_{ge}).$ (15)
The reader might wonder why the expectation value of the spin along $z$ depends on the coherences and not only on the difference in populations or why the $x$-component depends on the population. The reason is the inclusion of mixed states as a consequence of the magnetic field being tilted from the $z$ axis. Mathematically, Eq. 13 arises from changing the basis set from the classical up and down states of the Pauli matrix to the eigenstate basis set of the quantum impurity (QI) since ${\hat{S}}_{i}$ has to be expressed in the same basis as the reduced density matrix, meaning that we cannot simply use the standard Pauli matrices representation. This can be interpreted as a projection of the populations and coherences defined in the spin’s quantization axis onto the Cartesian coordinates. Equivalently, one could rotate the cartesian coordinates to the magnetic field direction, chaging the basis of the reduced density matrix instead, a procedure done in the literature [34, 21, 24, 18, 66] We emphasize that $\langle\mathbf{S}\rangle$ describes a quantum-statistical average and every component can adopt any number between -1/2 and 1/2.
Introducing Eq. 13 into Eq. 12 allows us to obtain an electric current equation that clearly shows the dependence on $\theta$. To simplify the result, we place ourselves in the long time where the populations are constant, i.e. $\rho_{ee(gg)}=P_{e(g)}$. Additionally, we consider that the linear driving term dominates ($A_{L}^{2}\sim 0$) so that the coherences are continuously Larmor precessing only, i.e. $\rho_{ge}(t)=\rho_{ge,-1}e^{i\omega t}+\rho_{ge,1}e^{-i\omega t}$, where the ±$\pm 1$ in the coherences refers to its Floquet number [36]. This yields:
$I_{L}$$\approx$ $\left(\mbox{Re}\left\{{\cal I}_{L}^{+}(\Delta_{2e})\right\}-\mbox{Re}\left\{{\cal I}_{L}^{-}(\Delta_{0e})\right\}\right)$
-- $P_{L}\left[(P_{e}-P_{g})\cos\theta-A_{L}(\rho_{ge,1}+\rho_{eg,-1})\sin\theta\ \right]$
$\times$ $\left(\mbox{Re}\left\{{\cal I}_{L}^{+}(\Delta_{2e})\right\}+\mbox{Re}\left\{{\cal I}_{L}^{-}(\Delta_{0e})\right\}\right)\implies$
$I_{L}$$\approx$ $I_{0}-P_{L}\left[\Delta P_{eg}\cos\theta\right.\left.-A_{L}(\rho_{ge,1}+\rho_{eg,-1})\sin\theta\ \right]I_{0}^{\prime},$
where the new definitions $I_{0},\ I_{0}^{\prime}$, which contain a non-trivial DC bias dependence, see Eqs. 9 and 10, and $\Delta P_{eg}$ can be deduced by comparing terms.
Equation LABEL:homo_current_final shows the clear importance of the spin polarization: if $P_{L}=0$, the electric current will have a constant unpolarized value due to electrons or holes hopping from the left to right (or vice versa) passing through the QI. However, this electric current does not depend on the populations and coherences of the QI, making the current measurement incapable of detecting changes in the QI state and, by extension, it cannot cause any spin flip or resonant behavior even if there is a driving applied. This happens because only a non-zero spin polarization imbues an imbalance in the populations for the up and down states and a change in the tunneling magnetoresistance. If $P_{L}\neq 0$, a polarized current emerges and $\langle S_{z}\rangle$ starts to play a role. On resonance and in the long time limit, the population of both singly occupied levels will be $1/2$, assuming finite but small coherences. In such a situation, $\langle S_{z}\rangle\rightarrow 0$. On the other hand, if $P_{L}\neq 0$ but $A_{L}=0$, the QI will have a net spin polarization $\langle S_{z}\rangle\neq 0$. However, the states cannot exchange population since the off coherences are zero, leading to no resonance at the Larmor frequency. Another key insight from Eq. LABEL:homo_current_final is the dependence on the homodyne angle. Just as in the absence of driving, when $\theta=0$, no exchange occurs between the populations since $\sin\theta=0$. In contrast, at $\theta=\pi/2$, the electric current becomes independent of the populations, resulting in an ESR signal governed purely by the coherences. Thus, Eq. LABEL:homo_current_final clearly demonstrates that homodyne detection arises from both the misalignment between the tip and impurity spin and the presence of AC driving.
II.4 Mapping onto a spin problem
Following Ref. [38], we now aim to map the QME (Eq. 5) onto a system of equations for the spin. This will become particularly useful for building some physical intuition on the definitions of the Rabi rate and the renormalization of the resonance energy. Using Eqs. 1314 and 15, the system of equations for the spin components and the population of the transient states is [^4] Equations LABEL:spin_equations assume a negligible Zeeman energy in the rates compared to the ionization and charge energy, $\varepsilon$ and $\varepsilon+U$. Since these energy thresholds are $\sim 100$ meV or more while $\Delta_{eg}=g\mu_{B}B<0.1$ meV, this is valid in the STM-ESR context  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. This allows us to write $\sum_{\alpha}P_{\alpha}\left[\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}-\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0})\right]\sin\theta=\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg}(t))$, which further simplifies the spin system of equations; see Appendix D.:
$\hbar\langle\dot{S}_{z}\rangle$ $=\Delta_{ge}\langle{S}_{y}\rangle\sin\theta-\hbar\langle{S}_{z}\rangle/\tau_{\mathrm{rel}}+\hbar\langle\dot{S}_{z}\rangle_{\mathrm{acc}}$ (17)
$\hbar\langle\dot{S}_{x}\rangle$ $=\left[2\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg}(t))/\sin\theta-\Delta_{ge}\cos\theta\right]\langle{S}_{y}\rangle-\hbar\langle{S}_{x}\rangle/\tau_{\mathrm{rel}}$
$\hbar\langle\dot{S}_{y}\rangle$ $=-2\mbox{Im}(\Gamma_{ev,vg}(t))\langle S_{x}\rangle/\sin\theta-\hbar\langle{S}_{y}\rangle/\tau_{\mathrm{rel}}-\Delta_{ge}\left(\langle S_{z}\rangle\sin\theta-\langle S_{x}\rangle\cos\theta\right)$
$\hbar\dot{\rho}_{0}$ $=2\mbox{Re}(\Gamma_{g0,0g}^{P_{\alpha}=0}(t))(1-\rho_{0}-\rho_{2})+4\sum_{\alpha}\mbox{Re}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}(t))P_{\alpha}\langle S_{z}\rangle/\hbar-4\mbox{Re}(\Gamma_{g0,0g}^{P_{\alpha}=0}(t))\rho_{0}$
$\hbar\dot{\rho}_{2}$ $=2\mbox{Re}(\Gamma_{g2,2g}^{P_{\alpha}=0}(t))(1-\rho_{0}-\rho_{2})-4\sum_{\alpha}\mbox{Re}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}(t))P_{\alpha}\langle S_{z}\rangle/\hbar-4\mbox{Re}(\Gamma_{g2,2g}^{P_{\alpha}=0}(t))\rho_{2}.$
The notation $\Gamma_{lj,uv}^{P_{\alpha}=0}(t)$ indicates zero spin polarization for the rates in Eq. 6. It is important to emphasize that the rates $\Gamma_{ev,vg}(t)$, that “flip” the excited state to the ground one or vice versa, are proportional to the spin polarization and $\sin\theta$, making them a consequence of both the lead spin polarization and the homodyne angle, i.e. those rates are exactly zero when $P_{L}=0$ and $\theta=0$. Thus $2\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg}(t))/\sin\theta$ does not depend on $\theta$.
Equations LABEL:spin_equations introduces the spin relaxation rate $1/\tau_{\mathrm{rel}}$ and the non-equilibrium spin accumulation $\langle\dot{S}_{z}\rangle_{\mathrm{acc}}$, parallel to the spin polarization as defined in Ref. [38],
$1/\tau_{\mathrm{rel}}$ $=\sum_{v=0,2}\mbox{Re}(\Gamma_{gv,vg}^{P_{\alpha}=0}(t)+\Gamma_{ev,ve}^{P_{\alpha}=0}(t))/\hbar$ (18)
$\langle\dot{S}_{z}\rangle_{\mathrm{acc}}$ $=\sum_{\alpha}P_{\alpha}\Bigg{[}2\mbox{Re}(\Gamma_{0g,g0,\alpha}^{P_{\alpha}=0}(t))\rho_{0}-2\mbox{Re}(\Gamma_{2g,g2,\alpha}^{P_{\alpha}=0}(t))\rho_{2}$ (19)
$+(\mbox{Re}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}(t)-\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}(t)))(1-\rho_{0}-\rho_{2})\Bigg{]},$
The spin accumulation term in Eq. 19 depends on the populations of transient and single occupied states of the QI, and describes the spin accumulation via tunneling to and from spin-polarized leads, which renders it proportional to the current. In short, the spin accumulation term is a source term that builds up the spin polarization on the QI. The spin relaxation rate Eq. 18 describes the decay of every spin component towards the steady state by electrons with a specific spin tunneling out of the QI, a process with $v=0$, or by tunneling in of a second electron with an opposite spin state, forming a doubly occupied state ($v=2$). The spin accumulation and relaxation govern the spin dynamics of the QI if the spin polarization and magnetic field are parallel, $\theta=0$. However, this is not the complete picture of the dynamics since the angle $\theta$ enters Eqs. LABEL:spin_equations. This orientation dependence can be encompassed in the components of the following rotational term:
$\langle\mathbf{\dot{S}}\rangle_{\mathrm{rot}}$ $=\mathbf{S}\times\sum_{\alpha}g\mu_{B}\mathbf{B_{\alpha}}(t)/\hbar$ (20)
A direct comparison with Eq. LABEL:spin_equations shows that:
$\mathbf{B_{\alpha}}(t)=\left(B\sin\theta,0,\frac{2\sum_{v}\mbox{Im}(\Gamma_{ev,vg,\alpha}(t))}{g\mu_{B}\sin\theta}+B\cos\theta\right).$ (21)
In this expression, we identify the so-called exchange magnetic field as
$g\mu_{B}\mathbf{B_{exch}}(t)={2\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg}(t))/\sin\theta}\ \mathbf{z}$ (22)
$=$ $-\sum_{\alpha}\frac{\gamma_{\alpha}(t)}{2\pi}P_{\alpha}\int^{\infty}_{-\infty}d\epsilon\left(\frac{\epsilon-\varepsilon}{(\epsilon-\varepsilon)^{2}+(\hbar/\tau_{c})^{2}}f_{\alpha}^{-}(\epsilon)\right.$
$+$ $\left.\frac{\epsilon-\varepsilon-U}{(\epsilon-\varepsilon-U)^{2}+(\hbar/\tau_{c})^{2}}f_{\alpha}^{+}(\epsilon)\right)\ \mathbf{z},$
[…]
In the CB regime, where the spin accumulation is small, the dynamics of the spin are dominated by Eq. 20 as long as $\theta\neq 0,\pi$. These two extreme cases are interesting though since they imply a simple relaxation of the $z$ component of the spin to the steady without any evolution of the $x$ and $y$ components, see Eq. LABEL:spin_equations. In the absence of driving, the exchange field will generate a spin splitting of the QI, a Zeeman splitting, even at $B=0$. When the driving is turned on, both the rotation and accumulation term will exhibit time-dependent behavior that will yield FLT and STT resonance signals in the frequency domain respectively.
III Hopping-driven SAIM: Rabi rate, $T_{2}$ times and energy dressing
In the following, we will discuss features of our model, continuing to use a spin-1/2 impurity. This approach results in shorter and more manageable expressions, leading to a more intuitive understanding of the underlying physics. We emphasize that while higher-spin systems can significantly alter the equations presented here by introducing additional singly and transient states, the central conclusions regarding the driving mechanism (FLT and STT) and the transport (above or below the energy thresholds) will remain valid.
III.1 Rabi rate
In previous studies, the Rabi frequency was introduced as a mathematical consequence of the model, conveying that the modulation of the tunneling barrier generates a driving of reasonable amplitude [27, 35, 36]. However, this analysis does not provide an intuitive understanding of the actual driving effect on the spin, thus it is useful to cast the driving mechanism into a more intuitive physical picture.
Here, we want to address the definition of the Rabi frequency by looking at the system of equations for the spin components for each of the two driving regimes considered in this work: FLT and STT. For the FLT-driven regime, which is presence in the CB regime, the simulations show the typical behavior of a spin under a time-dependent magnetic field, i.e. the spin performs Rabi oscillations. This points to the exchange field as the dominant source of driving. Contrary, in the STT-driven regime, presence above the thresholds that limit the CB regime, the spin only decays to a polarized steady state while performing Larmor oscillations in the plane perpendicular to the quantization axis of the spin. This is the result of the accumulation term being a dominant influence on the spin dynamics.
III.1.1 Rabi rate in the FLT-driven regime
There are two essential ingredients for the Rabi rate to emerge naturally in the QME: the spin polarization of one of the leads and the driving amplitude. We already showed the consequence of introducing polarization on the rates by mapping the QME onto a spin problem in Eqs. LABEL:spin_equations: the spin accumulation and rotation terms emerge as convenient descriptors for the spin dynamics. In the CB regime, the spin accumulation is rather small and can be partly discarded. Thus, only the rotation and relaxation terms strongly influence the time evolution of the spin.
When the exchange field and the spin are forming an angle $\theta\neq 0,\pi$, the exchange field creates a torque-like force on the initial spin that, together with the spin relaxation, will evolve the QI to the steady state. If this FLT produced by the exchange field is time-dependent, it is easy to conclude that it will induce a Rabi oscillation during this relaxation process, $\pi/2$ radians dephased respect to the driving term [24]. The strength of this Rabi rate will be given by the component of the exchange field torquing the spin, which is:
$\Omega_{\mathrm{FLT}}\propto B_{\mathrm{exch}}\cos(\pi/2-\theta).$
Looking at Eq. 22 we can easily identify
$2\hbar\Omega_{\mathrm{FLT}}=g\mu_{B}B_{\mathrm{exch}}\sin(\theta)=2\sum_{v=0,2}\mbox{Im}(\Gamma_{gv,ve}).$ (23)
This simple equation of the Rabi rate shows its physical meaning clearly as the projection of the exchange field on the perpendicular axis to the spin quantization in the plane where the spin and the exchange field lie. A comparison between QME and the optical Bloch equations also provides the same identification of the Rabi rate as the imaginary part of the rate associated to a spin-flip since only the imaginary part can provide Rabi oscillations. As a clear consequence of this definition, the Rabi rate is maximized at $\pi/2+m\pi$ with $m$ being an integer. However, this does not directly translate into a higher ESR signal amplitude because of the homodyne detection mechanism: the ESR signal measured by the electric current is a measurement of the $z$ component of the spin, as we will show in the simulation results.
Note that the exchange field includes up to second harmonic generation since
$\gamma_{\alpha}(t)=\gamma_{\alpha}^{0}\left[1+\frac{A_{\alpha}^{2}}{2}+2A_{\alpha}\cos(\omega t)+\frac{A_{\alpha}^{2}}{2}\cos(2\omega t)\right],$ (24)
but it could be expanded to higher harmonics if one considers the full Bessel function expression [67, 68] or a more complex oscillating behavior in 24. Following this, we can define the Rabi rates per harmonic $n=0,1,2$ as:
$\hbar\Omega^{0\omega}_{\mathrm{FLT}}$ $=$ $\frac{1}{2}g\mu_{B}B_{\mathrm{exch}}^{z}(A_{L}=0)\left(1+A_{L}^{2}/2\right)\sin\theta$
$\hbar\Omega^{1\omega}_{\mathrm{FLT}}$ $=$ $g\mu_{B}B_{\mathrm{exch}}^{z}(A_{L}=0)A_{L}\sin\theta$
$\hbar\Omega^{2\omega}_{\mathrm{FLT}}$ $=$ $\frac{1}{4}g\mu_{B}B_{\mathrm{exch}}^{z}(A_{L}=0)A_{L}^{2}\sin\theta$
Therefore, the exchange field provides zero-order and higher harmonic signals directly, which is in line with recent experiments on a strongly driven pentacene molecule spin-1/2 system [24, 69]. The $n$-th harmonic signal appears at a frequency of $f_{1\omega}/n$, with $f_{1\omega}$ being the first harmonic resonance frequency and $n>0$ the harmonic number. In this way, the second harmonic signal provides an additional resonance at $f_{1\omega}/2$ in the frequency spectrum. This can be easily verified since this 2th harmonic signal will emerge when $22\hbar\omega=4\pi\hbar f_{2\omega}=\Delta_{eg}\implies f_{2\omega}=\Delta_{eg}/2\hbar=f_{1\omega}/$ 2. If the frequency is fixed whilst sweeping the magnetic field, the $n$-th resonance magnetic field will roughly follow the relation $nB_{1\omega}$ with $B_{1\omega}$ the magnetic field of the first resonance [24, 69]. The zero-harmonic feature is a special case, and it is found at zero frequency in the frequency domain and near zero in the magnetic field domain.
In Ref. [24], the authors observed a peak near zero magnetic field with an intensity comparable to or slightly higher than that of the first harmonic peak. Notably, this zero-field peak appears even in the absence of driving. At first glance, based on the equations above, this result could be explained by the 0-th harmonic resonance, as both emerge at zero driving and require only a spin-polarized tip. However, the results in [24] were obtained in the STT-driven regime, making them more consistent with explanations based on spin bottlenecking [70] or the Hanle effect [43, 42]. Measurements in the FLT-driven regime of this phenomenon are still lacking, but they could offer coherent control of a single spin at zero AC driving, using only DC voltages.
Coming back to the equations, one can show that the resonance for the zero-harmonic signal occurs when $B=-B_{\mathrm{exch}}\cos\theta$ when the frequency is fixed. This suggests that this signal position can be used as a measurement of the exchange field. From our model’s perspective, the origin of this zero-field feature arises from spin-polarized electron precession in a magnetic field generated by the exchange interaction with the spin polarized tip. However, in the absence of driving, no ESR signal is expected for any harmonic when sweeping over the frequency for a fixed magnetic field, even at zero frequency, since the $\cos\omega t$ term vanishes in the rates.
III.1.2 Rabi rate in the STT-driven regime
When increasing the DC bias $eV_{\mathrm{DC}}$ beyond the energy thresholds, $eV_{\mathrm{DC}}<\varepsilon$ and $eV_{\mathrm{DC}}>\varepsilon+U$, we are outside the CB regime and the electric current through the system drastically increases at constant tip-atom distance. In this regime, the role of the exchange field is suppressed by the dominance of the spin accumulation term and the rotation term starts to decrease logarithmically. This does not imply the lack of ESR signal when the QI is driven in the STT regime. In fact, we can define the Rabi rates in this regime since the rates in the spin accumulation term (Eq. 19) can be written as the real part of a rate linked to a Rabi process by applying the following relations:
$\hbar\Omega_{\mathrm{STT}}(t)$ $=$ $\sum_{v=0,2}\mbox{Re}(\Gamma_{gv,ve}(t))$
$=$ $\sin\theta\sum_{\alpha}P_{\alpha}\left[\mbox{Re}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}(t)-\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}(t))\right]$
$\hbar\Omega_{\mathrm{STT}}^{v=0,2}(t)$ $=$ $\mbox{Re}(\Gamma_{vg,ev}(t))$
$=$ ±$\pm\sum_{\alpha}\mbox{Re}(\Gamma_{vg,gv,\alpha}^{P_{\alpha}=0}(t))P_{\alpha}\sin\theta,$
where the $+$ sign applies to $v=0$ while the -- sign is for $v=2$. The definitions of the Rabi rates above are easily confirmed by comparing our QME with the optical Bloch equation and considering that the Rabi rate is a complex number with a real and an imaginary part. If driving is applied, the Rabi rates above will modulate the spin accumulation, providing an ESR purely governed by the coherences [36]. On resonance, this forces the spin to undergo Larmor precession, in phase with the driving term, while it evolves to the steady state.
It is important to highlight that, in this regime, there is an inherent competition between the external magnetic field and the coupling to the spin polarized electrode, which determines where the spin will accumulate. If $B\gg\gamma_{L}^{0}$, the spin accumulation term is too weak to align the spin with the $z$-axis, and instead, the spin accumulates in the magnetic field orientation performing Larmor precession in the plane perpendicular to the spin’s quantization axis. Conversely, if $B\ll\gamma_{L}^{0}$, the spin accumulation term dominates, forcing the spin to align with the $z$-axis. This mechanism underlies the Hanle effect outside the CB regime [43, 42]. However, in both cases discussed, the sign of the spin accumulation determines whether the QI is spin polarized positively or negatively, disregarding the magnetic field orientation. In this work, we consider values of the coupling and magnetic field that faithfully reflect experimental conditions, by obtaining FLT Rabi frequencies in agreement with the experimental measurements, placing ourself in the $B\gg\gamma_{L}^{0}$ regime.
A key consequence of Larmor precession occurring in the plane perpendicular to the spin’s quantization axis is that the populations remain unchanged even at resonance, a result that persists in the long-time limit. This implies the absence of Rabi oscillations since, essentially, the spin is being forcedly accumulated in the direction set by the external magnetic field. Moreover, the ESR signal does not follow the homodyne detection [34, 21], leading to a fully symmetric resonance [36]. Another key feature of this regime is that Rabi rates are proportional to the electric current, exhibiting anharmonic behavior, since both the current and the Rabi rate depend on the coupling to the tip. As a result, the STT-driven regime naturally gives rise to higher harmonic resonances [69].
III.1.3 Analytical form of the Rabi
We could encompass both the FLT and STT Rabi rates in a complex rate term as $\hbar\Omega(t)=\hbar(\Omega_{\mathrm{STT}}(t)+i\Omega_{\mathrm{FLT}}(t))=\sum_{v=0,2}\Gamma_{gv,ve}(t)$. Every rate is calculated numerically in the wide band limit, which mandates a cutoff in the integral. If the cutoff energy, $E_{c}$, is much larger than any other energy in the QI, i.e. $E_{c}\gg|\varepsilon|,|\varepsilon+U|$, the rates will converge to a value independent of the cutoff [38, 63, 42], which is the situation in all our calculations. In this case, at extremely low temperatures, $k_{B}T_{\alpha}\ll|\varepsilon-\mu_{\alpha}|,|\varepsilon+U-\mu_{\alpha}|$, we can assume that the Fermi distribution function behaves as a step function:
$f_{\alpha}^{+}(\epsilon)$$\approx$ $1-\Theta(\epsilon-\mu_{\alpha})$
$f_{\alpha}^{-}(\epsilon)$$\approx$ $\Theta(\epsilon-\mu_{\alpha}),$
with $\Theta(x)$ being the Heaviside step function. This approximation allows us to write the simplest [^5] An analytical expression for the imaginary part of the Rabi rate at finite temperatures can be expressed using digamma functions [71] but we are interested in the low temperature scenario. analytical form of the Rabi rate for a spin-$1/2$ system as:
$\hbar\Omega(t)$ $=$ $-\sum_{\alpha}\frac{1}{4\pi}\gamma_{\alpha}^{0}(1+A_{\alpha}\cos(\omega t))^{2}P_{\alpha}\sin\theta$ (25)
$\times$ $\left[\arctan\left(\frac{\mu_{\alpha}-\varepsilon-U}{\hbar/\tau_{c}}\right)+\arctan\left(\frac{\mu_{\alpha}-\varepsilon}{\hbar/\tau_{c}}\right)\right.$
$+$ $\left.\frac{i}{2}\log\left(\frac{(\mu_{\alpha}-\varepsilon-U)^{2}+(\hbar/\tau_{c})^{2}}{(\mu_{\alpha}-\varepsilon)^{2}+(\hbar/\tau_{c})^{2}}\right)\right].$
In the limit $\hbar/\tau_{c}\ll|\varepsilon|,\varepsilon+U$, the real part of the Rabi frequency becomes close to zero in the CB regime, which results in nearly zero current as well. In contrast, outside of the CB regime, the sum of the $\arctan$ terms gives $\mbox{sign}(V_{\mathrm{DC}})\pi=\mbox{sign}(\mu_{L})\pi$, which depends on the sign of the chemical potential of the polarized electrode. On the other side, the imaginary part of Eq. 25 does not vanish in the CB regime, indicating its independence from the current, as we mentioned before.
III.2 Spin relaxation time and quality factor $\Omega T_{2}$
Let us move on to the spin relaxation Eq. 18. Since our model does not consider any pure dephasing mechanism and we have not included cotunneling processes, the time independent part of $\tau_{\mathrm{rel}}$ is actually the coherence time $T_{2}$, which is purely lifetime-limited so that $T_{2}=2T_{1}$. Assuming low temperature again, one arrives at an analytical expression for $T_{2}$:
$\frac{1}{T_{2}}$ $=$ $\frac{1}{h}\sum_{\alpha}\gamma_{\alpha}^{0}\left(1+\frac{A_{\alpha}^{2}}{2}\right)$ (26)
$\times$ $\int^{\infty}_{-\infty}d\epsilon\left[f_{\alpha}^{-}(\epsilon)\frac{\hbar/\tau_{c}}{(\epsilon-\varepsilon)^{2}+(\hbar/\tau_{c})^{2}}\right.$
$+$ $\left.f_{\alpha}^{+}(\epsilon)\frac{\hbar/\tau_{c}}{(\epsilon-\varepsilon-U)^{2}+(\hbar/\tau_{c})^{2}}\right]$
$\approx$ $\frac{1}{h}\sum_{\alpha}\gamma_{\alpha}^{0}\left(1+\frac{A_{\alpha}^{2}}{2}\right)\left[\pi-\arctan\left(\frac{\mu_{\alpha}-\varepsilon}{\hbar/\tau_{c}}\right)\right.$
$+$ $\left.\arctan\left(\frac{\mu_{\alpha}-\varepsilon-U}{\hbar/\tau_{c}}\right)\right],$
An important feature of Eq. 26 is that it saturates rapidly as we increase $|V_{\mathrm{DC}}|$ outside of the CB regime to
$\frac{1}{T_{2}}=\frac{1}{2\hbar}\gamma_{L}^{0}\left(1+\frac{A_{L}^{2}}{2}\right),$ (27)
where the right chemical potential is grounded and its contribution can be discarded compared to the left one for $eV_{\mathrm{DC}}=\mu_{L}<\varepsilon$ and $eV_{\mathrm{DC}}>\varepsilon+U$. Equation 27 does not depend on the induced broadening $\hbar/\tau_{c}$, since the DC bias is above the energy thresholds and is essentially controlled by the time independent left coupling $\gamma^{0}_{L}$, which is where the driving is applied and the tip-atom distance is encoded.
In the CB regime, however, the coherence times depend on the induced broadening, and a finite value of $\hbar/\tau_{c}$ prevents $T_{2}$ from diverging. This feature is a clear limitation of the model, as it was mentioned before in the introduction, and going beyond it requires extending to higher orders in perturbation theory of the tunneling amplitude $t_{\alpha}$, which will lead to cotunneling [39, 27]. Future extensions of this model will aim to circumvent such limitation, nonetheless, we set $\hbar/\tau_{c}\ll|\varepsilon|,\varepsilon+U$ so its influence on Eq. 26 is minimized. In this way, we can explore the behavior of $T_{2}$ in the CB regime although, at low DC voltage, we expect low current in line with previous studies [38, 62, 39, 63, 42] as the cotunneling regime becomes more significant in the deep CB regime.
The product $\Omega T_{2}$ per harmonic is the quality factor of a qubit, giving the number of Rabi oscillation the spin performs on resonance before reaching the steady state. We focus on the Rabi frequency related to $1\omega$, the first harmonic, under reasonable ESR conditions with $eV_{\mathrm{DC}}=\mu_{L}$ and the spin polarization and driving are only applied to the left lead. When the QI is driven under STT, the real part of the Rabi rate dominates, and we obtain:
$\Omega_{\mathrm{STT}}^{1\omega}T_{2}\approx-\frac{\mathrm{sign}(V_{\mathrm{DC}})\sin\theta}{2}\frac{A_{L}P_{L}}{\left(1+\frac{A_{L}^{2}}{2}\right)}.$ (28)
Equation 28 is proportional to the ESR amplitude and already suggests an optimal value for the driving. Additionally, it indicates that the ESR signal is independent of the couplings to the leads. If the driving is increased substantially, Eq. 28 will tend toward zero, and the ESR signal will diminish as well. Specifically, the driving value that maximizes the first harmonic signal is $\sqrt{2}$, yielding $\Omega_{\mathrm{STT}}^{1\omega}T_{2}=-\mathrm{sign}(V_{\mathrm{DC}})\sin\theta P_{L}\sqrt{2}/4$. Even in the ideal scenario of a fully polarized electrode and a magnetic field perpendicular to the exchange field, $|\Omega_{\mathrm{STT}}^{1\omega}T_{2}|=\sqrt{2}/4\approx 0.35$, meaning that in the STT-driven regime, we are unable to create a robust spin-qubit, as it will complete barely 1/3 of a Rabi oscillation before fully decaying to the steady state. This does not improve further if we consider the surviving Rabi rate of the FLT regime instead. Therefore, the spin is always overdamped in this scenario, and the ESR driven by STT is mainly useful for polarizing the QI as an initialization step, although the ESR signal can still be reasonably strong under suitable conditions [35, 36].
On the other hand, the quality factor in the FLT-driven regime, $\Omega_{\mathrm{FLT}}^{1\omega}T_{2}$, offers more flexibility to achieve larger values because, first, the Rabi rate does not vanish in the CB regime, and second, a reasonable $T_{2}$ can be maintained in this regime while keeping the induced broadening $\hbar/\tau_{c}$ low. Considering again that the DC bias applied to the tip and the sample is grounded, we arrive at:
$\Omega_{\mathrm{FLT}}^{1\omega}T_{2}\approx\frac{-i\gamma_{L}^{0}A_{L}P_{L}\sin\theta\log\left(\frac{(eV_{\mathrm{DC}}+\varepsilon+U)^{2}+(\hbar/\tau_{c})^{2}}{(eV_{\mathrm{DC}}+\varepsilon)^{2}+(\hbar/\tau_{c})^{2}}\right)/2}{\gamma_{L}^{0}\left(1+\frac{A_{L}^{2}}{2}\right)\left[\pi-\arctan\left(\frac{eV_{\mathrm{DC}}-\varepsilon}{\hbar/\tau_{c}}\right)+\arctan\left(\frac{eV_{\mathrm{DC}}-\varepsilon-U}{\hbar/\tau_{c}}\right)\right]+\gamma_{R}^{0}\left[\pi+\arctan\left(\frac{\varepsilon}{\hbar/\tau_{c}}\right)-\arctan\left(\frac{\varepsilon+U}{\hbar/\tau_{c}}\right)\right]}.$ (29)
Equation 29 depends on both left and right couplings, unlike Eq. 27, since now $\varepsilon<eV_{\mathrm{DC}}<\varepsilon+U$ and both left and right terms are in a similar order of magnitude. It is also required to retain the broadening terms to avoid divergences as we are considering low temperatures. Increasing the coupling of the unpolarized electrode to values where $\gamma_{R}\gg\gamma_{L}$ suppresses the ESR signal since $\Omega_{\mathrm{FLT}}^{1\omega}T_{2}\rightarrow 0$, while increasing $\gamma_{L}$ causes the ESR amplitude to reach a saturation point that becomes independent of the couplings. The former implies influence from the substrate, indicating that extremely strong coupling to it leads to no ESR. Equation 29 depends on the DC voltage and enables further optimization of the qubit’s quality factor beyond the limit of 1/3 of the STT-driven regime, as we will detail in the results section. Both extreme situations between the couplings yield an electric current independent of the spin polarization, suggesting an ideal factor between $\gamma_{R}$ and $\gamma_{L}$ that maximizes the ESR signal in the FLT-driven regime.
It is worth noting that cotunneling influences the ESR signal through the decoherence time $T_{2}$, which yields, for example, unexpected DC bias behavior with the ESR signal that our model cannot reproduce, as shown in [30]. Nonetheless, the strength of our model lies in the correct characterization of the Rabi rates through the exchange field and the spin accumulation, meaning that the driving is correctly accounted for and naturally given by the model itself once an AC electric field is considered.
Outside of the CB regime, Eq. 23 does not fully vanish, but it logarithmically tends toward zero. However, the $T_{2}$ drops much faster, making Eq. 29 extremely small for any driven regime, preventing any quantum control. Thus, both FLT and STT regimes will give a measurable ESR signal in the electric current, though their origins and utilities differ significantly as we will show in the results section: STT-ESR provides a strong spin polarization of the QI, whereas for the FLT-ESR, the QI can be reliably used as a spin-qubit.
III.3 Energy dressing
Returning to the exchange field, we observed that the FLT Rabi rate can be interpreted as the projection of the exchange field onto the axis perpendicular to the spin in the plane containing both the exchange field and the spin itself. However, we initially ignored the projection of this field along the spin axis. On the quantization axis of the spin the Hamiltonian is diagonal, meaning that the projection of the exchange on this axis will change the energy of the states, usually referred to as dressing of the energy. This projection provides a Zeeman-like splitting even in the absence of an external magnetic field, as long as the exchange field and spin are not perpendicular. When an external magnetic field is applied, the energy shift will modify the resonance frequency of the spectrum, shifting it relative to the bare resonance frequency given by $f_{0}=\Delta_{eg}/h=g\mu_{B}B/h$. Following this geometric interpretation, we define the energy (or frequency) shift as:
$\Delta E(t)$ $=$ $h\Delta f(t)=g\mu_{B}B_{\mathrm{exch}}(t)\cos\theta$
$=$ $2\sum_{\alpha}P_{\alpha}\left[\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}(t)-\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}(t))\right]\cos\theta,$
The time-dependent component in Eq. LABEL:energy_dress does not affect the spin dynamics, as it aligns with the quantization axis, though we include it here for completeness.
By neglecting the Zeeman energy relative to the charge energies, we obtain a useful relation between the Rabi rate and the energy shift:
$h\Delta f(t)\sin\theta=2\hbar\Omega_{\mathrm{FLT}}(t)\cos\theta$ (31)
This relation was used to derive Eqs. LABEL:spin_equations. Equation 31 provides a straightforward method to effectively measure the Rabi frequency, given the knowledge of the frequency shift and the angle between the tip and the external magnetic field. Another important detail about the energy dressing is how it affects zero-harmonic and higher-harmonic signals. Writing Eq. 5 in Floquet components, as in [35, 36], shows that in general the resonance frequency per harmonic is $hf_{n\omega}=(\Delta_{eg}+h\Delta f)/n$ which makes the energy shift dependent on the harmonic (or Floquet number) $n$. Interestingly, if we move to the perspective of the magnetic field being swept at a fixed frequency, the resonance magnetic field per harmonic will be $B_{n\omega}=h(nf-\Delta f)/g\mu_{B}$. This implies that the shifting in the magnetic field domain is the same per harmonic since $\Delta f$ is not been altered by $n$. As a consequence, the zero-harmonic feature in the frequency domain appears at zero frequency, while in the magnetic field domain, it does at $-h\Delta f/g\mu_{B}=-B_{\mathrm{exch}}\cos\theta$. This offset can be used to measure the exchange field indirectly, a proposal in line with other studies [42].
IV Results
In this section, we examine key points from the previous discussion of the theory through numerical analysis, highlighting the differences between ESR signals of the first harmonic only driven by STT and FLT. Model parameters were chosen to ensure a strong ESR signal from an $S=1/2$ system in the FLT-driven regime while keeping the system weakly coupled to the leads under driving, as per similar requirements in [36]. We summarize the parameters employed in our simulations next.
The left lead represents the spin-polarized tip with a spin polarization of $P_{L}=0.8$ along with the findings in [30]. The coupling strengths are set to $\gamma_{R}^{0}=3\times\gamma_{L}^{0}=15\ \mu$eV for all simulations and the temperature is equal to $T=1$ K for both the left and right leads. In the context of ESR-STM experiments, this implies that the tip-sample distance is constant and the background current will change accordingly with the DC bias. The DC bias drop is applied to the tip so $eV_{DC}=\mu_{L}$ since we align the chemical potential of the right lead with the Fermi energy, which defines the zero energy. The singly occupied energy is $\varepsilon=-100$ meV, and the Coulomb repulsion is $U=250$ meV. These values offer a realistic enough exploration of both tunneling regimes not so far off from DFT and others predictions [25, 72]. We have increased the intrinsic width of the impurity states to $\hbar/\tau_{c}=0.25$meV, compared to the one used in [36], allowing exploration of the CB regime while maintaining $\hbar/\tau_{c}\ll|\varepsilon|,\ \varepsilon+U$. The driving amplitude is set to $A_{L}=50\%$, applied only to the polarized lead, consistent with the experimental setup. Finally, the Zeeman splitting provides a bare resonance frequency of approximately $f_{0}=16.8$ GHz, with an isotropic $g$-factor of 2. The energy dressing will modify the resonance position to a value $f_{0}^{\prime}=\Delta f+f_{0}$, see Eq. LABEL:energy_dress, which is taken into account to perform on resonance simulations. Additional parameters are specified in figure captions.
IV.1 Spin accumulation vs. DC bias without AC driving
[…]
The effect of spin accumulation persists even in the CB regime, albeit weaker, vanishing completely at zero voltage. There is, however, a competition between the relaxation time, which ranges from 0.2 to 50 ns (Fig. 7a), and $\hbar/\dot{S}_{\mathrm{acc}}$, leading to different results depending on the DC bias. In the CB regime, $\hbar/\dot{S}_{\mathrm{acc}}\gg\tau_{\mathrm{rel}}$, so the spin evolution is dominated by the relaxation term, as the accumulation term acts extremely slowly on the QI. This opens the possibility of an FLT-ESR signal since the time dependent exchange field will have ”time” to torque the spin coherently, producing Rabi oscillations. On the other hand, above the energy thresholds outside of the CB regime, both the relaxation and spin accumulation terms are of the same order of magnitude, forcing the QI to be spin-accumulated in the quantization axis before any coherent evolution occurs. Therefore, AC driving that leads to a time-dependent spin accumulation will provide an incoherent STT-ESR signal, as we will see in the following sections.
IV.2 ESR in the STT-driven regime
[…]
Figure 3: ESR outside of the CB regime; out of plane magnetic field. a) Bloch sphere representation of the spin components in the lab frame for $\theta=\pi/18$ and positive $eV_{\mathrm{DC}}/|\varepsilon|=2$. The spin is initialized thermally at $\langle\mathbf{S}\rangle_{\mathrm{thermal}}=(-0.033,0,-0.188)$ (red arrow). The positive DC bias causes the spin to flip to its spin polarization in the steady state as a consequence the spin accumulation term. At the same time, the spin continuously performs Larmor oscillations even in the steady state. The Bloch sphere axes are reduced for clarity. b) Time-evolution of the current from the thermal state, reaching the steady state after 2 ns. c) Corresponding ESR electric current signal in the STT-driven regime in the frequency domain. The resonance frequency is slightly different to the bare one at 16.8 GHz being now $\Delta f+f_{0}=17.07$ GHz due to the energy dressing.
Figure 4: ESR outside of the CB regime; in-plane magnetic field. a) Bloch sphere representation of the spin components in the lab frame for $\theta=\pi/2$ and $eV_{\mathrm{DC}}/|\varepsilon|=2$. The spin is thermally initialized at $\langle\mathbf{S}\rangle_{\mathrm{thermal}}=(-0.19,0,0)$ (red arrow). Here, the spin accumulation forces alignment with the quantization axis of the spin in the steady state, leading to continuous Larmor oscillations in the $zy$-plane. b) Time-evolution of the current from the thermal situation. c) ESR electric current signal in the STT regime for the case of a) in the frequency domain. Since in-plane initialization leads to not dressing, the resonance occurs at the bare one $f_{0}=16.8$ GHz.
Fig. 3b illustrates the spin dynamics, characterized by fast Larmor oscillations that persist even in the long-time limit. These oscillations occur in a plane perpendicular to the quantization axis, Fig. 3a, defined by the magnetic field, and arise from the time-dependent spin accumulation. This accumulation, proportional to the polarized current, attempts to torque the accumulated spin toward the spin polarization axis, which in result makes the crossed product $\mathbf{S}\times\mathbf{B}$ non-zero, forcing the spin to torque with an intensity proportional to $\sin\theta$. In this regime, the populations remain frozen, and the Larmor oscillations only modify the coherences, making the ESR current fully governed by them [36].
Fig. 3c shows the corresponding frequency behavior. Although the ESR signal is significant, the background current is three orders of magnitude higher, resulting in an extremely low ESR signal relative to the background. This is expected in the STT-driven regime, as reflected in the low-quality factor $\Omega_{\mathrm{STT}}T_{2}$. The signal exhibits an asymmetric Fano profile due to the exchange field and the residual presence of homodyne detection. However, it becomes fully symmetric when $eV_{\mathrm{DC}}\gg\varepsilon+U$, where the exchange field vanishes (see Eq. 25).
For the in-plane configuration, Fig. 4a, the strongest Larmor oscillations and ESR signals (Figs. 4b and 4c)) are achieved. The Larmor precession now occurs in the $zy$-plane, again perpendicular to the magnetic field and the quantization axis. If the STT-driven regime would follow the homodyne detection previously reported [34, 21], the ESR signal would be minimal or zero since the spin and tip spin polarization are perpendicular, and the ESR signal is proportional to their scalar product. However, in this regime, this limitation does not hold as the spin will be driven towards the $z$ axis by the spin accumulation. After this initial push, the spin feels the torque from the external magnetic field, from the cross product $\langle\mathbf{S}\rangle\times\mathbf{B}$. This torque is maximized for in plane configurations, producing a stronger ESR signal compared to the out of plane situation. This interaction allows the initial oscillatory evolution of the spin observed both in Fig. 4a and Fig. 3a. Shortly thereafter, the spin settles into oscillations within the $zy$ plane in Fig. 4a, driven by the time-dependent spin accumulation and the torque exerted by the external magnetic field.
The role of the exchange field in this scenario is analogous to the out-of-plane case; however, the ESR signal in Fig. 4c is nearly symmetrical. This symmetry arises because, when the spin is initialized along a Cartesian axis, $\langle S_{z}\rangle$ in Eq. 13 depends solely on either the populations or the decoherences, but not on a combination of both. Consequently, no Fano profile emerges. Specifically, for an in-plane magnetic field, the $z$ component of the spin depends only on the real part of the coherences and the $x$ one gives the populations. Importantly, with an initial in-plane configuration, the spin cannot polarize because the spin accumulation term is fully perpendicular to the external magnetic field. As a result, the system tends to equalize populations into a superposition state at the center of the Bloch sphere if no driving is applied. Note that in the in-plane configuration, the populations, measured in the $x$ axis, also remain unperturbed in the long-time limit, even at resonance, as observed in the out-of-plane situation Fig. 3.
The reader may notice that the ESR signal in both cases (Figs. 3c and 4c) are dips (a reduction in the current on-resonance) which is a direct result of the STT-driven regime at positive bias. In this regime, the coherences peak on resonance while remaining zero off resonance. According to Eq. 12, this behavior produces a dip in the signal. Conversely, under a negative DC bias ($eV_{\mathrm{DC}}<\varepsilon$), the current becomes negative and the ESR signal will flip from a dip to a peak. However, the ESR signal cannot flip by merely changing the sign of the spin polarization because an additional sign change occurs in the $z$ component of the spin, effectively preserving the overall response. Higher order tunneling processes could induce additional flipping of the signal though.
IV.3 ESR in the FLT-driven regime
We now shift our focus to the Coulomb blockade (CB) regime, where FLT-driven resonance dominates. This regime favors quantum-coherent manipulation of the spin over the spin polarization of the impurity spin since the spin accumulation term is significantly minimized. This reduces the impact of the completely decoherent polarized current on the impurity, enabling the spin to be used as a qubit. FLT driven regime is basically equivalent to conventional ESR where a time dependent magnetic field induces resonant transition. As highlighted in the theoretical discussion, this allows for quality factors $\Omega_{\mathrm{FLT}}T_{2}$ to reach their upper limits, constrained only by the ability to enhance $T_{2}$ for a fixed Rabi rate. Notably, this enhancement is strongly dependent on the DC bias, making it the critical factor in achieving high coherence for spin manipulation.
[…]
Figure 5: ESR in the Coulomb Blockade regime; out of plane magnetic field. a) Bloch sphere representation of the spin components in the rotating frame for a magnetic field tilted by $\theta=\pi/18$ and a bias voltage of $eV_{\mathrm{DC}}/|\varepsilon|=1.275$, where the maximum quality factor $\Omega T_{2}\approx 6$ is achieved. The spin is thermally initialized at $\langle\mathbf{S}\rangle_{\text{thermal}}=(-0.033,0,-0.188)$ and experiences a torque due to the time dependent exchange field. This torque causes the spin to precess in a plane with the normal vector $\mathbf{n}=(-\sin(4\pi/9),0,\cos(4\pi/9))$. b) Time evolution of the current showing up to six Rabi oscillation cycles before decaying to the steady state in the rotating frame. c) ESR signal, DC contribution, featuring a peak with a subtle Fano profile. The resonance frequency is slightly tuned with respect to the bare one to be $\Delta f+f_{0}=17.146$ GHz due to the energy dressing.
Figure 6: ESR in the Coulomb Blockade regime; in plane magnetic field. a) Bloch sphere representation of the spin components in the rotating frame for a magnetic field aligned with the plane ($\theta=\pi/2$) and a bias voltage of $eV_{\mathrm{DC}}/|\varepsilon|=1.275$, where the maximum quality factor $\Omega T_{2}\approx 36$ is achieved. The spin is thermally initialized at $\langle\mathbf{S}\rangle_{\text{thermal}}=(-0.19,0,0)$ and undergoes torque to the exchange field, leading to precession in a plane with the normal vector $\mathbf{n}=(0,0,1)$. b) Time evolution of the current, showing approximately 36 Rabi oscillations before decaying to the steady state in the rotating frame, closely matching the z-component of the spin, not the $x$-component. c) ESR signal, DC contribution, displaying a dip without Fano profile in the frequency domain. Since in-plane initialization leads to not dressing, the resonance frequency is $16.8$ GHz.
The current in the rotating frame, as shown in Fig. 5b, directly reflects the Rabi oscillations of the spin, as it corresponds to the $z$-component of the spin. Notably, the DC contribution here is significantly smaller than that in Fig. 3b, as we are in the CB regime. This reduced DC contribution allows for a relatively long coherence time $T_{2}$ of approximately 35 ns without diminishing the contribution of the exchange field. An interesting aspect is observed in the ESR signal (Fig. 5c), which now exhibits a sharp peak instead of the dips seen earlier. This occurs because the ESR signal depends not only on the coherences but also on the populations of the spin states when the spin Rabi oscillates. On resonance, the populations tend toward a 50/50 distribution if the driving force is sufficiently strong, as expected in classical ESR simulations. At positive bias and spin polarization, the off-resonance population in the long-time limit leads to $\langle S_{z}\rangle>0$, product of, albeit small, spin accumulation, and consequently, the ESR signal on resonance manifests as a peak, see Eq. 12. Conversely, at negative bias, the signal flips to a dip, with an inverted Fano profile.
We now turn to the time evolution of an in-plane initialized spin, Fig. 6a. The Rabi rate $\Omega_{\mathrm{FLT}}$ is maximized at $\theta=\pi/2$ radians, where the projection of the exchange field perpendicular to the spin is greatest, effectively maximizing the torque exerted by the exchange field. A notable feature in this simulation is the spin being pulled toward the positive $z$-axis of the Bloch sphere. This phenomenon can be attributed to the a non-negligible spin accumulation, which is expected given the DC bias’s proximity to the threshold $\varepsilon+U$. This residual spin accumulation explains a non-zero ESR signal even in the perpendicular configuration of the tip and spin—behavior, where typically no signal is found due to the homodyne detection [34].
Since the current measures the $z$-component of the spin, it does not reflect the Rabi oscillations as clearly as in the out-of-plane case, as confirmed in Fig. 6b. However, the current’s maxima and minima in the rotating frame still correspond to the spin being in the ground and excited states, respectively, demonstrating controllability of the spin. It is worth noting that while our theory lacks a fully developed cotunneling framework, it captures key aspects of the detection mechanism, particularly homodyne detection. This is evident from the intensity minimum of the ESR signal in Fig. 6c, despite the higher quality factor $\Omega_{\mathrm{FLT}}T_{2}$ in the in-plane configuration compared to the out-of-plane case Fig. 5c. Thus, we conclude that while our model has limitations, it successfully incorporates the homodyne detection, and it aligns with experimental results, indicating that out-of-plane magnetic fields yield stronger ESR signals than in-plane configurations.
If we examine the spin dynamics on the Bloch sphere in the lab frame instead of the rotating frame, the spin exhibits an additional Larmor oscillation perpendicular to the plane in which the spin undergoes Rabi oscillations. This arises from the cross product between the spin and the external magnetic field.
IV.4 Behavior of $\Omega$ and $T_{2}$ under DC bias
Figure 7: Rabi rate and coherence time behavior under DC bias a Coherence time $T_{2}$ as a function of the DC bias. A maximum of roughly 50 ns is achieved at the electron-hole symmetry point (vertical red dash-dotted line). The thresholds at $\varepsilon$ and $\varepsilon+U$ are marked as vertical blue dashed lines, after which the $T_{2}$ time drops to 0.23 ns. b) Rabi frequency behavior in both the FLT (solid line) and STT (dotted line) driven regimes as a function of DC bias. Both Rabi rates follow the equations from the theory section. (c) Figure of merit $\Omega T_{2}$, which determines the number of Rabi cycles and the efficiency of the ESR signal if the homodyne detection is not playing a role. Even in the absence of clear Rabi oscillations due to a small $T_{2}$, the product $\Omega T_{2}$ can still be large enough to produce a significant ESR signal.
In the following, we explore the behavior of the Rabi frequency $\Omega$ and coherence time $T_{2}$ as functions of the easily adjustable experimental parameter $eV_{\mathrm{DC}}$ for the out-of-plane situation studied in Figs. 3 and 5. The result for $T_{2}$ is shown in Fig. 7a while the one for the Rabi rates is depicted in Fig. 7b. We separate the contributions from the STT and FLT driven regimes in Fig. 7b, as these evolve quite differently under varying DC bias. The behavior of $T_{2}$ is relatively straightforward to understand based on the previous discussion. In the low-bias CB regime, the current through the impurity is not a full one-electron process. Since scattering between the transport electrons and the impurity is the primary source of decoherence in this model, the coherence time is maximized. Therefore, the CB regime effectively protects the impurity spin from the decoherence induced by the polarized current. As we approach the charging thresholds $\varepsilon$ and $\varepsilon+U$ getting out of the CB regime, $T_{2}$ rapidly decreases. Due to the single-orbital nature of our model, $T_{2}$ then reaches a constant and very small value once the DC bias $V_{\mathrm{DC}}$ exceeds these thresholds. This demonstrates that the choice of $V_{\mathrm{DC}}$ is critical for achieving favorable conditions for quantum control.
Fig. 7b illustrates the behavior of the Rabi rate in both the STT and FLT driven regimes. In the CB regime, the FLT Rabi rate dominates over the SST Rabi rates, following a logarithmic behavior, as indicated by the imaginary part of Eq. 25. For low biases, $\Omega_{\mathrm{FLT}}$ exhibits a linear dependence that does not vanish at zero DC bias, which explains the presence of an ESR signal at $V_{\mathrm{DC}}=0$ meV, even though the signal intensity is low due to the minimal current passing through the impurity. As the bias approaches the CB regime threshold, $\Omega_{\mathrm{FLT}}$ becomes highly nonlinear, reaching a maximum (minimum) at $\varepsilon+U$ ($\varepsilon$). According to Eq. 25, $\Omega_{\mathrm{FLT}}$ would theoretically diverge at $\varepsilon+U$ and $\varepsilon$ for $\hbar/\tau_{c}\ll\varepsilon+U,|\varepsilon|$, but finite temperature acts as a natural cutoff, preventing this divergence. Above the energy thresholds, the SST Rabi rate emerges strongly, leading to an incoherent ESR signal, as shown in Fig. 3 and 4. The sign of the Rabi rates in each driven regime determines the rotation direction: a negative sign induces clockwise rotation in the Bloch sphere, while a positive sign results in counterclockwise rotation.
At the electron-hole symmetry point, located at $eV_{\mathrm{DC}}=\varepsilon+U/2=25$ meV, the additional energy required to add or remove an electron is equal, balancing the electron and hole processes and causing the exchange field, $\Omega_{\mathrm{FLT}}$, to vanish, resulting in zero ESR signal. Notably, at this symmetry point, the coherence time $T_{2}$ reaches its maximum value, indicating that the impurity spin is most effectively shielded from the influence of the decoherent polarized current when the electron and hole processes cancel out each other. It is important to note that the electron-hole symmetry point may be absent if we do not apply the wide-band limit approximation in the density of states of the electrodes.
Figure 7c shows the figure of merit $\Omega T_{2}$, indicating a maximum of roughly 6 Rabi cycles before the steady state is reached for the FLT Rabi rate, as clearly depicted in Fig. 5. This underscores the importance of optimizing the DC bias to operate in the CB regime for achieving the best possible conditions for quantum control. Contrary, the $\Omega_{\mathrm{STT}}$ behavior only emerges outside the CB regime when the spin accumulation dominates the dynamics as Fig. 7b shows. Therefore, the product $\Omega_{\mathrm{STT}}T_{2}$ does not have a large value for any DC bias, as we inferred from Eq. 28. In the STT-driven regime, the spin is weakly driven, making this regime useless in terms of quantum control, but it can be utilized to polarize the QI to a high degree as we already showed in Fig. 3.
IV.5 Full DC voltage mapping and energy dressing
In this section, we examine the ESR signal response under varying DC bias for a fixed polar angle of $\theta=\pi/18$. The results are presented in Fig. 8, where a 2D color map highlights the normalized ESR signal relative to the background (baseline off-resonance) current, preserving the characteristic peak and dip behaviors. The efficiency of the ESR signal closely follows the figure of merit $\Omega_{\mathrm{FLT}}T_{2}$ shown in Fig. 7b, demonstrating low efficiency outside the CB regime beyond the energy thresholds $\varepsilon$ and $\varepsilon+U$. As the DC bias approaches zero, the background current diminishes, albeit at a different rate compared to $\Delta I$. This results in a non-zero ESR signal at zero bias, a phenomenon reported previously [19, 30]. This observation aligns with the earlier discussion that $\Omega_{\mathrm{FLT}}$ does not vanish at zero bias but rather at the electron-hole symmetry point, $eV_{\mathrm{DC}}/|\varepsilon|=1+U/2|\varepsilon|=0.25$. Indeed, there the signal completely disappears, as predicted by the exchange field formula, Eq. 23, and observed in Fig. 7.
Figure 8: Shifting of the ESR resonance position with DC bias. The ESR signal intensity, normalized to the background current, is shown as a color map, revealing peaks at positive DC biases and dips at negative DC biases. As the voltage approaches the thresholds $\varepsilon$ and $\varepsilon+U$, the ESR efficiency decreases substantially. The observed frequency shift agrees with the prediction of Eq. LABEL:energy_dress and deviates by up to approximately 8% (about 1.3 GHz) from the bare resonance frequency $f_{0}=g\mu_{B}B/h=16.8$ GHz.
An intriguing phenomenon observed in Fig. 8 is the shifting of the ESR resonance position. We emphasize that in these simulations, the tip-sample distance is kept constant, therefore, the couplings remain unchanged while sweeping the DC bias. This dressing of the resonance energy arises from the projection of the exchange field onto the quantization axis of the spin and only vanishes at the electron-hole symmetry point or when the spin is initialized exactly perpendicular to the exchange field. In these cases, the resonance occurs at the bare resonance frequency $f_{0}=g\mu_{B}B/h=16.8$ GHz. This frequency shift has been consistently reported in the literature  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] , with various explanations proposed. Recently, it was suggested to be a result of a change of the $g$-factor by the electric field of the DC bias [19]. However, the net effect of the DC bias on the magnetic moment found in experiments [25], supported by DFT simulations, is orders of magnitude too small and rules out $g$-factor variation as the leading cause. There are also considerations regarding Ref. [22], which rules out the $g$-factor as a contributing factor. A $g$-modulation would discard the utility of the Fe atoms to drive the single-spin 1/2 Ti atom, contradicting experimental evidence, a point we discuss further in Ref. [73]. This leaves the exchange field as the most plausible mechanism for this phenomenon, though minor contributions from other sources cannot be entirely ruled out.
The observable shift in the resonance frequency was used as a direct experimental method to determine the impurity parameters $\varepsilon$ and $\varepsilon+U$ [25], aligning well with DFT simulations [72]. Moreover, the shift depicted in Fig. 8 is clearly proportional to $\Omega_{\mathrm{FLT}}$ shown in Fig. 7a. Thus, measuring the resonance shift offers a practical way to obtain the Rabi rate. This prediction could be validated using alternative Rabi rate estimation methods, such as the one presented in [23].
IV.6 Vector field dependence of the Rabi rate and ESR signal
The polar angle $\theta$ formed between the external magnetic field and tip spin polarization appears naturally in all quantities of interest and is closely related to measure an efficient ESR signal, as we observed by comparing Figs. 5c and 6c. As such, it is important to understand the behavior of our model with respect to the polar angle. Furthermore, since the static magnetic field angle is a parameter easily adjusted in some experiments, confirming the agreement we have between model and experiment will provide validity of the driving field emerging from the modulation of the exchange magnetic field and the hopping.
Figure 9: Example of the ESR signal behavior on resonance, calculating the total amplitude $\Delta I_{p}=\sqrt{\Delta I_{s}^{2}+\Delta I_{a}^{2}}$, at negative (dotted line) and positive (solid line) DC bias as a function of the polar angle $\theta$ formed between the magnetic field and the tip spin polarization. The maximum ESR signal occurs at a predominantly out-of-plane angle, as a result of homodyne detection. A sizable signal for an in-plane angle arises due to the spin accumulation being noticeable for these DC biases.
As an example of the angle behavior we have Fig. 9, where we show the total ESR amplitude $\Delta I_{p}=\sqrt{\Delta I_{s}^{2}+\Delta I_{a}^{2}}$ for the optimal DC bias values that maximizes the quality factor $\Omega_{\mathrm{FLT}}T_{2}$, see Fig. 7c. $\Delta I_{s}$ denotes the symmetric part of the ESR signal, while $\Delta I_{a}$ provides the asymmetric one. Both are extracted from fitting the raw ESR signal to a symmetric and an antisymmetric Lorentzian center on the resonance frequency (which includes the frequency shift or energy dressing).
At $\theta=0,\pi$, the ESR signal completely vanishes as the driving field and the spin are parallel. This is not found experimentally due to the tip likely forming an angle $\theta^{\prime}$ with the plane containing the spin. In other words, in general, the tip spin polarization does not lie exactly on $z$ in the experiments, as it does in the model. This can be accounted for in the simulations by tilting the plane in which the magnetic field lies by an angle $\theta^{\prime}$, relative to the spin polarization in $z$.
Figure 10: Full DC and vector field dependence. a) 2D color map of the ESR total amplitude $\Delta I_{p}=\sqrt{\Delta I_{s}^{2}+\Delta I_{a}^{2}}$, normalized to the background current keeping its sign, as a function of the external field angle and DC bias, highlighting regions of high signal efficiency. $\Delta I_{s}$ gives the symmetric part of the ESR signal, while $\Delta I_{a}$ provides the asymmetric one. At $\theta=0,\pi$ and along the electron-hole symmetry lines, no ESR signal is observed. The white horizontal dashed lines correspond to the plots shown in Fig. 9. b) Figure of merit $\Omega_{\mathrm{FLT}}T_{2}$ as a function of the polar angle $\theta$ and DC bias, showing significant variation within the Coulomb blockade (CB) regime, proportional to $\sin\theta$, and a rapid drop-off outside of it. The red dotted line indicates the electron-hole symmetry line at $eV_{\mathrm{DC}}/|\varepsilon|=1+U/2|\varepsilon|=0.25$.
Outside of those special polar angles, both negative and positive DC bias in Fig. 9 show a similar behavior for the ESR amplitude that follows roughly the homodyne detection, as the largest ESR amplitude is found predominately out of plane ($\theta\sim 10$ degrees), which matches experimental evidence [34, 21]. However, at 90 degrees, we encounter a minimum of the ESR amplitude which goes against the conventional ESR-STM measurement. Here, the homodyne detection tells us that the amplitude should be zero for in-plane angle as the ESR signal is proportional to the scalar product of the spin and the tip direction. As we already mentioned, we identify this discrepancy as a limitation of the model at capturing the ESR signal in the FLT-driven regime, which includes a small spin accumulation contribution towards the $z$ axis, providing a non-zero ESR signal. A more robust theory that fully integrates the sequential tunneling and cotunneling transport is needed to obtain the full picture, the former being responsible for creating the exchange field. It is worth noting that, since we are plotting the ESR amplitude $\Delta I_{p}$, we do not visualize the flipping of the signal happening when the angle is sweeping from an out-of-plane value to an in-plane one, going through a pure fano-profile ESR signal.
Turning our sight to Fig. 10a, we observe the ESR amplitude dependence versus both the DC bias and the polar angle. The behavior with the voltage is similar to the one presented in Fig. 8 and independent of the angle, but without visualizing the resonance shifting since we are plotting the ESR amplitude on resonance. Instead, if we focus on the angle dependence, one can see again that the ESR signal will vanish in the case of parallel ($\theta=0^{\circ}$) and anti-parallel ($\theta=180^{\circ}$) orientations. As we pointed out in Fig. 9, the ESR amplitude has a local minimum for a perpendicular alignment ($\theta=\pm 90^{\circ}$) for any DC bias due to the spin accumulation, however, the lower the absolute value of the voltage is, the smaller the ESR amplitude is at the perpendicular alignment. This is because the spin accumulation is reduced even further the lower the DC bias is, see Fig. 2, even though the quality factor $\Omega_{\mathrm{FLT}}T_{2}$ maximizes at $\theta=90$ degrees as Fig. 10b indicates.
Additionally, Fig. 10b clearly shows the sinusoidal dependence of the Rabi rate on the polar angle, as experimentally reported in [21], since $T_{2}$ remains independent of the magnetic field orientation. The electron-hole symmetry line, $eV_{\mathrm{DC}}=\varepsilon+U/2$, is represented by a red dotted line in Fig. 10b, where both the Rabi rate and ESR signal vanish, consistent with Fig. 10a at $eV_{\mathrm{DC}}/|\varepsilon|=0.25$.
Outside the Coulomb blockade (CB) regime, the quality factor drops to extremely low values, following the decrease in $T_{2}$ observed in Fig. 7a. Near zero DC bias, the ESR amplitude relative to the background exhibits a local maximum or minimum, depending on whether $\theta\in(-90^{\circ},90^{\circ})$ or falls within the range $[90^{\circ},270^{\circ}]$, respectively. This ESR signal behavior aligns with the fact that the Rabi rate remains nonzero at this bias, as discussed earlier.
V Conclusions
In this work, we derived an intuitive physical picture of ESR in an STM, building on the solid mathematical aspects of non-equilibrium transport through a spin-polarized electrode under harmonic modulation of the bias. The key aspect of this model is the emergence of ESR driven by purely electric fields, which is experimentally realized in ESR-STM. We used a single-orbital Anderson impurity model with ionization and charging energies $\varepsilon,\varepsilon+U$, which is fairly general and can describe a wide variety of localized spin defects. Our model accounts for the driving amplitude in both the FLT and STT-driven regimes. However, it does not fully incorporate the complete transport picture, as cotunneling is not considered, leaving the door open for future extensions of this work.
As we mentioned, the model has two distinctively different transport regimes. When the DC bias lies within the charging window $\varepsilon,\varepsilon+U$ the transport occurs in the Coulomb blockade (CB) regime. Here, the impurity is singly occupied and we found that ESR in this region is driven by an effective magnetic exchange field $B_{\mathrm{exch}}$ that arises as a consequence of the spin polarization of the electrode $P_{\alpha}$ i.e. the QI is driven by a FLT and we obtain results that aligns with conventional ESR following the Homodyne detection and presenting Rabi oscillations. Its component parallel to the spin quantization axis leads to a renormalization of the resonance frequency $f_{0}$ which is logarithmically dependent on the applied bias relative to the charging energy of the SAIM. This shift can be used to spectroscopically probe the impurity levels $\varepsilon,\varepsilon+U$ and the Rabi rate proportional to the exchange field. We further show that the quality factor, the product of Rabi rate and coherence time $\Omega_{\mathrm{FLT}}T_{2}$, has an optimum value close to the charging energy thresholds. In general, in the CB regime, the impurity spin is an ideal candidate for quantum-coherent control.
When the bias crosses the charging thresholds, transport occurs outside the Coulomb blockade (CB) regime, allowing a full electron to hop in and out of the QI from the electrodes. Consequently, the current through the impurity increases, and $T_{2}$ decreases drastically. In this regime, an ESR signal can still be observed, with the driving mechanism now dominated by the spin polarized current, which torques the spin via the spin accumulation term and was referred as spin-transfer torque (STT). However, quantum-coherent control of the spin is not feasible due to rapid decoherence, as $\Omega_{\mathrm{STT}}T_{2}\sim 1/3$. Despite this, the QI can be strongly polarized along the spin’s quantization axis, provided it is not initialized perpendicular to the spin polarization. This offers an alternative method for initializing the spin in a preferred orientation, independent of temperature. Previously, STT was only studied using the semiclassical Landau–Lifshitz–Gilbert–Slonczewski equation that treats the magnetization as a classical vector. Thus, our model provides the first insight into quantum spin-transfer torque (qSTT).
As a final remark, our model is able to reproduce a range of experimental observation in qualitative and quantitative good agreement, such as the homodyne detection and resonance shifts, which solidifies the validity of our driving mechanism approach and can serve as a guide for numerical optimization of on-surface electron spin systems for quantum simulations, quantum sensing and quantum information processing as well as a leap forward in the pursuit of a more complete theory able to encompass every transport regime i.e. both sequential and cotunneling regimes without discarding the exchange field from the sequential one.
Acknowledgements.
We acknowledge fruitful discussions with A. Ardavan, A. J. Heinrich, C. Ast, C. P. Lutz, D.-J. Choi, D. Ruckert, D. Tomaszewski, F. Delgado, F. Donati, G. Burkard, G. Czap, H. T. Bui, J. Barnaś, J. König, J. C. Cuevas, L. Hoang-Anh, M. Kelai, M. Ternes, P. Kot, P. Busz, S.-H. Phark, S. Baumann, S. Loth, S. Stepanow, S. Kovarik and R. Schlitz during the process of writing this manuscript.
This work was supported by the Institute for Basic Science (IBS-R027-D1). Further financial support from projects RTI2018-097895-B-C44 and PID2021-127917NB-I00 funded by MCIN/AEI/10.13039/501100011033, QUAN-000021-01 funded by the Gipuzkoa Provincial Council and IT-1527-22 funded by the Basque Government is gratefully acknowledged. This work also received support from National Science Centre of Poland (Grant No. 2020/36/C/ST3/00539). Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Expectation value $\langle S_{i}\rangle$
In the main text, we introduced the general expression for the expectation values $\langle S_{i}\rangle=\mathrm{Tr}(\hat{\rho}\hat{S}_{i})$ with $i=x,y,z$. Here, we derive these expressions for a general magnetic field $\mathbf{B}=B(\sin\phi\sin\theta,\cos\phi\sin\theta,\cos\theta)$, where $\theta$ is the polar angle and $\phi$ the azimuthal angle.
A key consideration when computing the trace is ensuring that both operators are represented in the same basis. Otherwise, the expectation value would be incorrect. This is particularly relevant when considering an external magnetic field, as the quantization axis of the spin aligns with the magnetic field for an isotropic $g$-factor. Consequently, the density matrix and the Pauli matrices do not necessarily share the same basis, since the former is defined in the eigenstate basis, which is tilted by the magnetic field.
To circumvent this problem, one could change the basis set of either operator into the other one. In this work, we chose to perform a basis transformation of the standard Pauli matrices to the eigenvector basis set:
$\langle l|\hat{S}_{i}|j\rangle=\sum_{\sigma,\sigma^{\prime}}\langle l|\sigma\rangle\langle\sigma|\hat{S}_{i}|\sigma^{\prime}\rangle\langle\sigma^{\prime}|j\rangle,$
where $l,j$ are singly occupied eigenstates of the spin-1/2, while $\sigma,\sigma^{\prime}$ correspond to the spin-up and spin-down states and $S_{i}$ are the Pauli matrices. We focus exclusively on the singly occupied states, as the empty and doubly occupied states provide zero spin to the simplest QI. For a general magnetic field, the singly eigenstates $l$ and $j$ have the following expression:
$|g\rangle$ $=$ $\left(\sin\frac{\theta}{2},-e^{i\phi}\cos\frac{\theta}{2}\right)$
$|e\rangle$ $=$ $\left(\cos\frac{\theta}{2},e^{i\phi}\sin\frac{\theta}{2}\right),$ (32)
where $g$ and $e$ denote the ground and excite states of the QI respectively. Meanwhile, the standard up and down spin states basis set of the Pauli matrices are:
[…]
The density matrix is always expressed with the ground state as the first matrix element:
$\hat{\rho}=\begin{pmatrix}\rho_{gg}&\rho_{eg}\\ \rho_{ge}&\rho_{ee}\end{pmatrix}$
Thus, the first matrix element of the spin operator must correspond to the ground state. Computing every matrix element for the $z$-component leads to:
$\langle g|\hat{S}_{z}|g\rangle$ $=$ $\frac{\hbar}{2}\left(\langle g|\mkern-5.0mu\uparrow\rangle\langle\uparrow\mkern-5.0mu|g\rangle-\langle g|\mkern-5.0mu\downarrow\rangle\langle\downarrow\mkern-5.0mu|g\rangle\right)$
$=$ $\frac{\hbar}{2}\left(\sin^{2}\left(\frac{\theta}{2}\right)-\cos^{2}\left(\frac{\theta}{2}\right)\right)=-\frac{\hbar\cos\theta}{2}$
$\langle g|\hat{S}_{z}|e\rangle$ $=$ $\frac{\hbar\sin\theta}{2}\ \ ;\ \ \langle e|\hat{S}_{z}|g\rangle=\frac{\hbar\sin\theta}{2}$
$\langle e|\hat{S}_{z}|e\rangle$ $=$ $\frac{\hbar\cos\theta}{2}$
Thus we write
$\hat{S}_{z}=\frac{\hbar}{2}\begin{pmatrix}-\cos\theta&\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix},$
which is the general expression of the spin operator on z. This implies that:
$\hat{\rho}\hat{S}_{z}=\frac{\hbar}{2}\begin{pmatrix}-\rho_{gg}\cos\theta+\rho_{ge}\sin\theta&-\rho_{eg}\cos\theta+\rho_{ee}\sin\theta\\ \rho_{gg}\sin\theta+\rho_{ge}\cos\theta&\rho_{eg}\sin\theta+\rho_{ee}\cos\theta\end{pmatrix},$
arriving at
$\langle S_{z}\rangle$ $=$ $\frac{\hbar}{2}(\rho_{ee}-\rho_{gg})\cos\theta+\frac{\hbar}{2}(\rho_{ge}+\rho_{eg})\sin\theta$
$\underbrace{\approx}_{\theta\ll 1}$ $\frac{\hbar}{2}(\rho_{\uparrow\uparrow}-\rho_{\downarrow\downarrow}),$
which is the equation used in the main text. Only when $\theta$ is sufficiently small we recover the up and down notation, indicating that the spin is in the vertical axis of the Bloch sphere. Turning our attention to the $x$-component of the spin we obtain:
$\langle g|\hat{S}_{x}|g\rangle$ $=$ $\frac{\hbar}{2}\left(\langle g|\mkern-5.0mu\uparrow\rangle\langle\downarrow\mkern-5.0mu|g\rangle+\langle g|\mkern-5.0mu\downarrow\rangle\langle\uparrow\mkern-5.0mu|g\rangle\right)$
$=$ $-\frac{\hbar}{2}\left(e^{i\phi}+e^{-i\phi}\right)\sin\frac{\theta}{2}\cos\frac{\theta}{2}$
$=$ $-\frac{\hbar}{2}\cos\phi\sin\theta$
$\langle g|\hat{S}_{x}|e\rangle$ $=$ $-\frac{\hbar}{2}(\cos\theta\cos\phi-i\sin\phi)$
$\langle e|\hat{S}_{x}|g\rangle$ $=$ $-\frac{\hbar}{2}(\cos\theta\cos\phi+i\sin\phi)$
$\langle e|\hat{S}_{x}|e\rangle$ $=$ $\frac{\hbar}{2}\cos\phi\sin\theta$
Hence the transformed operator is
$\hat{S}_{x}=-\frac{\hbar}{2}\begin{pmatrix}\cos\phi\sin\theta&\cos\theta\cos\phi+i\sin\phi\\ \cos\theta\cos\phi-i\sin\phi&-\cos\phi\sin\theta\end{pmatrix},$
which leads to the expectation value
$\langle{S_{x}}\rangle$ $=$ $\frac{\hbar}{2}(\rho_{ee}-\rho_{gg})\sin\theta\cos\phi$
-- $\frac{\hbar}{2}(\rho_{ge}+\rho_{eg})\cos\theta\cos\phi-i\frac{\hbar}{2}(\rho_{ge}-\rho_{eg})\sin\phi.$
Finally, for the $y$-component we have
$\langle g|\hat{S}_{y}|g\rangle$ $=$ $\frac{\hbar}{2}\left(-i\langle g|\mkern-5.0mu\uparrow\rangle\langle\downarrow\mkern-5.0mu|g\rangle+i\langle g|\mkern-5.0mu\downarrow\rangle\langle\uparrow\mkern-5.0mu|g\rangle\right)$
$=$ $i\frac{\hbar}{2}\left(e^{i\phi}\sin\frac{\theta}{2}\cos\frac{\theta}{2}-e^{-i\phi}\sin\frac{\theta}{2}\cos\frac{\theta}{2}\right)$
$=$ $-\frac{\hbar}{2}\sin\phi\sin\theta$
$\langle g|\hat{S}_{y}|e\rangle$ $=$ $-\frac{\hbar}{2}(i\cos\phi+\cos\theta\sin\phi)$
$\langle e|\hat{S}_{y}|g\rangle$ $=$ $\frac{\hbar}{2}(i\cos\phi-\cos\theta\sin\phi)$
$\langle e|\hat{S}_{y}|e\rangle$ $=$ $\frac{\hbar}{2}\sin\phi\sin\theta$
Therefore,
$\hat{S}_{y}=-\frac{\hbar}{2}\begin{pmatrix}\sin\phi\sin\theta&-i\cos\phi+\cos\theta\sin\phi\\ i\cos\phi+\cos\theta\sin\phi&-\sin\phi\sin\theta\end{pmatrix},$
and we obtain
$\langle{S_{y}}\rangle$ $=$ $\frac{\hbar}{2}(\rho_{ee}-\rho_{gg})\sin\theta\sin\phi$
-- $\frac{\hbar}{2}(\rho_{ge}+\rho_{eg})\cos\theta\sin\phi+i\frac{\hbar}{2}(\rho_{ge}-\rho_{eg})\cos\phi.$
For simplicity, we have been considering the case of $\phi=0$ for the derivation of Eqs. LABEL:spin_equations and throughout this work, arriving at the expressions for the spin of x and y used in the main text. From the spin components, we can obtain the population difference as well as the real and imaginary parts of the coherences, which are defined along the spin quantization axis:
$\frac{\rho_{ee}-\rho_{gg}}{2}$ $=$ $\frac{\langle S_{z}\rangle\cos\theta+\langle S_{x}\rangle\sin\theta\cos\phi+\langle S_{y}\rangle\sin\theta\sin\phi}{\hbar}$
$\frac{\rho_{ge}+\rho_{eg}}{2}$ $=$ $\frac{\langle S_{z}\rangle\sin\theta-\langle S_{x}\rangle\cos\theta\cos\phi-\langle S_{y}\rangle\cos\theta\sin\phi}{\hbar}$
$\frac{\rho_{ge}-\rho_{eg}}{2}$ $=$ $\frac{i}{\hbar}\left(\langle S_{x}\rangle\sin\phi-\langle S_{y}\rangle\cos\phi\right).$ (33)
Appendix B Coefficients $\sum_{\sigma}\lambda_{lj,\sigma}\mu_{uv,\sigma}(1+2\sigma P_{\alpha}))$
These coefficients appear in the rates, and knowing how to evaluate them is essential for computing certain equations in this work. For the general eigenstates of a spin-1/2, Eqs. 32, the non-zero matrix elements between the ground/excited states and the transient ones are:
$\lambda_{0g,\uparrow}=\langle 0|0\rangle\langle\uparrow\mkern-5.0mu|g\rangle=\sin\frac{\theta}{2}$
$\lambda_{0e,\uparrow}=\langle 0|0\rangle\langle\uparrow\mkern-5.0mu|e\rangle=\cos\frac{\theta}{2}$
$\lambda_{0g,\downarrow}=\langle 0|0\rangle\langle\downarrow\mkern-5.0mu|g\rangle=-e^{i\phi}\cos\frac{\theta}{2}$
$\lambda_{0e,\downarrow}=\langle 0|0\rangle\langle\downarrow\mkern-5.0mu|e\rangle=e^{i\phi}\sin\frac{\theta}{2}$
$\lambda_{g2,\uparrow}=\langle g|\mkern-5.0mu\downarrow\rangle\langle 2|2\rangle=-e^{-i\phi}\cos\frac{\theta}{2}$
$\lambda_{e2,\uparrow}=\langle e|\mkern-5.0mu\downarrow\rangle\langle 2|2\rangle=e^{-i\phi}\sin\frac{\theta}{2}$
$\lambda_{g2,\downarrow}=\langle g|\mkern-5.0mu\uparrow\rangle\langle 2|2\rangle=\sin\frac{\theta}{2}$
$\lambda_{e2,\downarrow}=\langle e|\mkern-5.0mu\uparrow\rangle\langle 2|2\rangle=\cos\frac{\theta}{2}$
Taking into account that $\gamma_{\alpha\sigma}^{0}=\gamma_{\alpha}^{0}(1/2+\sigma P_{\alpha})$, we can compute all the relevant coefficients:
$\sum_{\sigma}\lambda_{0g,\sigma}\mu_{g0,\sigma}(1+2\sigma P_{\alpha})$
$=$ $(1+P_{\alpha})\sin^{2}\frac{\theta}{2}+(1-P_{\alpha})\cos^{2}\frac{\theta}{2}=1-P_{\alpha}\cos\theta$
$\sum_{\sigma}\lambda_{0e,\sigma}\mu_{e0,\sigma}(1+2\sigma P_{\alpha})$
$=$ $(1+P_{\alpha})\cos^{2}\frac{\theta}{2}+(1-P_{\alpha})\sin^{2}\frac{\theta}{2}=1+P_{\alpha}\cos\theta$
$\sum_{\sigma}\lambda_{g2,\sigma}\mu_{2g,\sigma}(1+2\sigma P_{\alpha})=1+P_{\alpha}\cos\theta$
$\sum_{\sigma}\lambda_{e2,\sigma}\mu_{2e,\sigma}(1+2\sigma P_{\alpha})=1-P_{\alpha}\cos\theta$
$\sum_{\sigma}\lambda_{0g,\sigma}\mu_{e0,\sigma}(1+2\sigma P_{\alpha})=(1+P_{\alpha})\sin\frac{\theta}{2}\cos\frac{\theta}{2}$
-- $(1-P_{\alpha})\cos\frac{\theta}{2}\sin\frac{\theta}{2}=P_{\alpha}\sin\theta$
$\sum_{\sigma}\lambda_{g2,\sigma}\mu_{2e,\sigma}(1+2\sigma P_{\alpha})=-P_{\alpha}\sin\theta.$
Appendix C Details on current expression for a spin 1/2
We will refer to results from the previous two appendices A and B for $\phi=0$, as in the main text. The current leaving the left lead is
$I_{L}(t)$ $=$ $\frac{2e}{\hbar}\left[\rho_{ee}(t)\mbox{Re}\left\{\Gamma_{e2,2e,L}^{+}(t)-\Gamma_{e0,0e,L}^{-}(t)\right\}\right.$ (34)
$+$ $\left.\rho_{gg}(t)\mbox{Re}\left\{\Gamma_{g2,2g,L}^{+}(t)-\Gamma_{g0,0g,L}^{-}(t)\right\}\right.$
$+$ $\mbox{Re}\{\rho_{ge}(t)\}\mbox{Re}\left\{\Gamma_{g2,2e,L}^{+}(t)-\Gamma_{g0,0e,L}^{-}(t)\right.$
$+$ $\left.\Gamma_{e2,2g,L}^{+}(t)-\Gamma_{e0,0g,L}^{-}(t)\right\}$
$+$ $\sum_{j=e,g}\left(\rho_{0}(t)\mbox{Re}\left\{\Gamma_{0jj0,L}^{+}\right\}-\rho_{2}(t)\mbox{Re}\left\{\Gamma_{2jj2,L}^{-}\right\}\right)$
$=$ $I_{L}^{2}(t)+I_{L}^{0}(t)=-I_{R}(t),$
where we have separated hole and electron contributions to the current and canceled out the ones that are zero due to being matrix elements $\lambda_{lj,\sigma}$ null. The major difference between the hole and electron rates comes from the first one containing $\varepsilon+U$, while the second one $\varepsilon$, but the spin dependence is similar and encoded in terms such as $\sum_{\sigma}\lambda_{lv\sigma}\lambda_{vj\sigma}^{*}(1+2\sigma P_{L})$, which we have already computed in the previous appendix. These expressions allow us to write:
$\Gamma_{ev,ve,L}^{\pm}(t)=\frac{1}{2}(1\mp P_{L}\cos\theta)\gamma_{L}(t){\cal I}_{L}^{\pm}(\Delta_{ve})$
$\Gamma_{gv,vg,L}^{\pm}(t)=\frac{1}{2}(1\pm P_{L}\cos\theta)\gamma_{L}(t){\cal I}_{L}^{\pm}(\Delta_{vg})$
$\Gamma_{ev,vg,L}^{\pm}(t)=\mp\frac{1}{2}P_{L}\sin\theta\gamma_{L}(t){\cal I}_{L}^{\pm}(\Delta_{vg})$
$\sum_{j=e,g}\Gamma_{0j,j0,L}^{+}(t)=\gamma_{L}(t){\cal I}_{L}^{+}(\Delta_{j0})$
$\sum_{j=e,g}\Gamma_{2j,j2,L}^{-}(t)=\gamma_{L}(t){\cal I}_{L}^{-}(\Delta_{j2})$
where, on the first three equations, the ones where the first subscript in the rates start with a singly occupied state, the $+$ sign signifies the electron-like transport, implying that $v=0$, while the -- sign applies to the hole-like process, where $v=2$. Additionally, we have introduced the time dependent non-polarized coupling $\gamma_{L}(t)=\left(1+A_{L}\cos(\omega t)\right)^{2}\gamma_{L}^{0}$. Since the magnetic field is in the GHz regime, the integrals ${\cal I}_{L}^{\pm}(\Delta_{vj})$ will not depend on the Zeeman energy but only on $\varepsilon$, $U$, and $\mu_{\alpha}$ as we assumed in the main text, so we have:
$I_{L}^{0}(t)$ $=$ $-\frac{e}{\hbar}\gamma_{L}^{0}\left(1+A_{L}\cos(\omega t)\right)^{2}\left(\mbox{Re}\left\{{\cal I}_{L}^{-}(\Delta_{0e})\right\}\right.$
$\times$ $\left[\rho_{ee}(t)+\rho_{gg}(t)+P_{L}\cos\theta(\rho_{ee}(t)-\rho_{gg}(t))\right.$
$+$ $\left.2\mbox{Re}\{\rho_{ge}(t)\}P_{L}\sin\theta\right]-\left.2\rho_{0}\mbox{Re}\left\{{\cal I}_{L}^{+}(\Delta_{e0})\right\}\right).$
Introducing Eq. 13 leads to
$I_{L}^{0}(t)$ $=$ $-\frac{e}{\hbar}\gamma_{L}(t)\left[(\rho_{1}+2P_{L}\langle S_{z}\rangle)\mbox{Re}\left\{{\cal I}_{L}^{-}(\Delta_{0e})\right\}\right.$
-- $\left.2\rho_{0}\mbox{Re}\left\{{\cal I}_{L}^{+}(\Delta_{e0})\right\}\right],$
and similarly
$I_{L}^{2}(t)$ $=$ $\frac{e}{\hbar}\gamma_{L}(t)\left[(\rho_{1}-2P_{L}\langle S_{z}\rangle)\mbox{Re}\left\{{\cal I}_{L}^{+}(\Delta_{2e})\right\}\right.$
-- $\left.2\rho_{2}\mbox{Re}\left\{{\cal I}_{L}^{-}(\Delta_{e2})\right\}\right],$
which together gives Eq. 12 when the population of the transient states, $\rho_{0}$ and $\rho_{2}$, are negligible and $\rho_{ee}(t)+\rho_{gg}(t)=\rho_{1}=1-\rho_{2}-\rho_{0}\approx 1$. We have employed Eq. 13. An analytical form of the current at low temperature is easy to obtain in terms of $\arctan$ functions.
Appendix D Insights on mapping the QME onto a spin problem
Here we aim to clarify some of the steps done for mapping the spin 1/2 version of Eqs. 5 into a spin system obtaining Eqs. LABEL:spin_equations. This means that we will derive a set of equations that depends solely on the spin components and the transient populations of the QI. To begin with, let us write the full set of equations for the populations and coherences for the simplest QI:
$\hbar\dot{\rho}_{gg}$ $=$ $-2\sum_{v=0,2}\mbox{Re}(\Gamma_{gv,vg}(t))\rho_{gg}-\sum_{v=0,2}\Gamma_{ev,vg}(t)\rho_{ge}-\sum_{v=0,2}\Gamma_{ev,vg}^{*}(t)\rho_{eg}+2\sum_{v=0,2}\mbox{Re}(\Gamma_{vg,gv}(t))\rho_{v}$
$\hbar\dot{\rho}_{ge}$ $=$ $-\sum_{v=0,2}\Gamma_{ev,vg}(t)\rho_{gg}+\left[i\Delta_{ge}-\sum_{v=0,2}\left(\Gamma_{ev,ve}(t)+\Gamma_{gv,vg}^{*}(t)\right)\right]\rho_{ge}-\sum_{v=0,2}\Gamma_{ev,vg}^{*}(t)\rho_{ee}+2\sum_{v=0,2}\mbox{Re}(\Gamma_{vg,ev}(t))\rho_{v}$
$\hbar\dot{\rho}_{ee}$ $=$ $-\sum_{v=0,2}\Gamma_{ev,vg}^{*}(t)\rho_{ge}-2\sum_{v=0,2}\mbox{Re}(\Gamma_{ev,ve}(t))\rho_{ee}-\sum_{v=0,2}\Gamma_{ev,vg}(t)\rho_{eg}+2\sum_{v=0,2}\mbox{Re}\Gamma_{ve,ev}(t)\rho_{v}$
$\hbar\dot{\rho}_{eg}$ $=$ $-\sum_{v=0,2}\Gamma_{ev,vg}^{*}(t)\rho_{gg}-\sum_{v=0,2}\Gamma_{ev,vg}(t)\rho_{ee}-\left[i\Delta_{ge}+\sum_{v=0,2}\left(\Gamma_{gv,vg}(t)+\Gamma_{ev,ve}^{*}(t)\right)\right]\rho_{eg}+2\sum_{v=0,2}\mbox{Re}(\Gamma_{vg,ev}(t))\rho_{v}$
$\hbar\dot{\rho}_{0}$ $=$ $2\mbox{Re}(\Gamma_{e0,0e}(t))\rho_{ee}+2\mbox{Re}(\Gamma_{g0,0g}(t))\rho_{gg}+2\mbox{Re}(\Gamma_{e0,0g}(t))(\rho_{ge}+\rho_{eg})-2\mbox{Re}(\Gamma_{0g,g0}(t)+\Gamma_{0e,e0}(t)))\rho_{0}$
$\hbar\dot{\rho}_{2}$ $=$ $2\mbox{Re}(\Gamma_{e2,2e}(t))\rho_{ee}+2\mbox{Re}(\Gamma_{g2,2g}(t))\rho_{gg}+2\mbox{Re}(\Gamma_{e2,2g}(t))(\rho_{ge}+\rho_{eg})-2\mbox{Re}(\Gamma_{2g,g2}(t)+\Gamma_{2e,e2}(t)))\rho_{2}.$
At this point, we need to make use of several relations for the rates that applies when both the magnetic field and the frequency of the time dependent field are small compared to the ionization and charge energies $\varepsilon$ and $\varepsilon+U$. Choosing $\pm$ depending on $v=0,2$ respectively we have that: $\mbox{Re}(\Gamma_{ev,ve})+\mbox{Re}(\Gamma_{gv,vg})=2\mbox{Re}(\Gamma_{ev,ve}^{P_{\alpha}=0}),$ $\mbox{Re}(\Gamma_{ev,ve})-\mbox{Re}(\Gamma_{gv,vg})=\pm 2\sum_{\alpha}P_{\alpha}\cos\theta\ \mbox{Re}(\Gamma_{gv,vg,\alpha}^{P_{\alpha}=0}),$ $\mbox{Im}(\Gamma_{ev,ve})+\mbox{Im}(\Gamma_{gv,vg})=2\mbox{Im}(\Gamma_{gv,vg}^{P_{\alpha}=0}),\mbox{Im}(\Gamma_{ev,ve})-\mbox{Im}(\Gamma_{gv,vg})=\pm 2\sum_{\alpha}P_{\alpha}\cos\theta\ \mbox{Im}(\Gamma_{ev,ve,\alpha}^{P_{\alpha}=0}),$ and $\Gamma_{ev,vg}=\Gamma_{gv,ve}$. In fact, in general, we can write: $\Gamma_{ev,ve}=\sum_{\alpha}\Gamma_{ev,ve,\alpha}^{P_{\alpha}=0}(1\pm P_{\alpha}\cos\theta)$ and $\Gamma_{gv,vg}=\sum_{\alpha}\Gamma_{gv,vg,\alpha}^{P_{\alpha}=0}(1\mp P_{\alpha}\cos\theta)$. With these relations and Eqs. 1314 and 15 we can write the spin system of equations:
$\langle\dot{S}_{z}\rangle$ $=$ ${\cos\theta}\left[-\sum_{v=0,2}\mbox{Re}(\Gamma_{gv,vg}^{P_{\alpha}=0})(\rho_{ee}-\rho_{gg})+i\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg})\rho_{ge}-i\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg})\rho_{ge}^{*}\right.$
$+$ $\left.\sum_{\alpha}P_{\alpha}\cos\theta\left(-\mbox{Re}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0})(\rho_{ee}+\rho_{gg})+\mbox{Re}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0})(\rho_{ee}+\rho_{gg})+2\mbox{Re}(\Gamma_{0g,g0,\alpha}^{P_{\alpha}=0})\rho_{0}-2\mbox{Re}(\Gamma_{2g,g2}^{P_{\alpha}=0})\rho_{2}\right)\right]$
$+$ ${\sin\theta}\left[-\sum_{v=0,2}\mbox{Re}(\Gamma_{ev,vg})(\rho_{gg}+\rho_{ee})+2\sum_{v=0,2}\mbox{Re}(\Gamma_{vg,ev})\rho_{v}\right.$
-- $\left.(\Delta_{ge}-2\sum_{\alpha}P_{\alpha}\cos\theta(\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0})-\mbox{Im}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}))\mbox{Im}(\rho_{ge})-2\sum_{v=0,2}\mbox{Re}(\Gamma_{gv,vg}^{P_{\alpha}=0})\mbox{Re}(\rho_{ge})\right],$
$\langle\dot{S}_{x}\rangle$ $=$ ${\sin\theta}\left[-\sum_{v=0,2}\mbox{Re}(\Gamma_{gv,vg}^{P_{\alpha}=0})(\rho_{ee}-\rho_{gg})+i\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg})\rho_{ge}-i\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg})\rho_{ge}^{*}\right.$
$+$ $\left.\sum_{\alpha}P_{\alpha}\cos\theta\left(-\mbox{Re}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0})(\rho_{ee}+\rho_{gg})+\mbox{Re}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0})(\rho_{ee}+\rho_{gg})+2\mbox{Re}(\Gamma_{0g,g0,\alpha}^{P_{\alpha}=0})\rho_{0}-2\mbox{Re}(\Gamma_{2g,g2,\alpha}^{P_{\alpha}=0})\rho_{2}\right)\right]$
-- ${\cos\theta}\left[-\sum_{v=0,2}\mbox{Re}(\Gamma_{ev,vg})(\rho_{gg}+\rho_{ee})+2\sum_{v=0,2}\mbox{Re}(\Gamma_{vg,ev})\rho_{v}\right.$
-- $\left.(\Delta_{ge}-2\sum_{\alpha}P_{\alpha}\cos\theta(\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0})-\mbox{Im}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}))\mbox{Im}(\rho_{ge})-2\sum_{v=0,2}\mbox{Re}(\Gamma_{gv,vg}^{P_{\alpha}=0})\mbox{Re}(\rho_{ge})\right],$
$\langle\dot{S}_{y}\rangle$ $=$ $\left[\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg})(\rho_{ee}-\rho_{gg})+(\Delta_{ge}-2\sum_{\alpha}P_{\alpha}\cos\theta(\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0})-\mbox{Im}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}))\mbox{Re}(\rho_{ge})\right.$
-- $\left.2\sum_{v=0,2}\mbox{Re}(\Gamma_{gv,vg}^{P_{\alpha}=0})\mbox{Im}(\rho_{ge})\right],$
which can be expressed independently of the singly occupied state populations and coherences by introducing $\hbar\mbox{Re}(\rho_{ge})=\langle S_{z}\rangle\sin\theta-\langle S_{x}\rangle\cos\theta$, $\langle S_{y}\rangle=-\hbar\mbox{Im}(\rho_{ge})$, and $\rho_{ee}-\rho_{gg}=2(\langle S_{z}\rangle\cos\theta+\langle S_{x}\rangle\sin\theta)/\hbar$, which follow from Eqs. 33 when $\phi=0$. This leads to:
$\langle\dot{S}_{z}\rangle$ $=$ $\left[\sum_{v=0,2}2{\cos\theta}\mbox{Im}(\Gamma_{ev,vg})+\sin\theta(\Delta_{ge}-2\sum_{\alpha}P_{\alpha}\cos\theta(\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0})-\mbox{Im}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}))\right]\langle S_{y}\rangle/\hbar-2\sum_{v=0,2}\mbox{Re}(\Gamma_{gv,vg}^{P_{\alpha}=0})\langle{S}_{z}\rangle/\hbar$
$+$ $\left.\sum_{\alpha}P_{\alpha}\cos^{2}\theta\left([\mbox{Re}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0})-\mbox{Re}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0})](1-\rho_{0}-\rho_{2})+2\mbox{Re}(\Gamma_{0g,g0,\alpha}^{P_{\alpha}=0})\rho_{0}-2\mbox{Re}(\Gamma_{2g,g2,\alpha}^{P_{\alpha}=0})\rho_{2}\right)\right.$
$+$ ${\sin\theta}\left[2\mbox{Re}(\Gamma_{g0,0g,e0})\rho_{0}+2\mbox{Re}(\Gamma_{2g,e2})\rho_{2}-(\mbox{Re}(\Gamma_{e0,0g})+\mbox{Re}(\Gamma_{e2,2g}))(1-\rho_{0}-\rho_{2})\right],$
$\langle\dot{S}_{x}\rangle$ $=$ $\left[\sum_{v=0,2}2{\sin\theta}\mbox{Im}(\Gamma_{ev,vg})-\cos\theta(\Delta_{ge}-2\sum_{\alpha}P_{\alpha}\cos\theta(\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0})-\mbox{Im}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}))\right]\langle S_{y}\rangle/\hbar-2\sum_{v=0,2}\mbox{Re}(\Gamma_{gv,vg}^{P_{\alpha}=0})\langle{S}_{x}\rangle/\hbar$
$+$ $\left.\sum_{\alpha}P_{\alpha}\cos\theta\sin\theta\left([\mbox{Re}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0})-\mbox{Re}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0})](1-\rho_{0}-\rho_{2})+2\mbox{Re}(\Gamma_{0g,g0,\alpha}^{P_{\alpha}=0})\rho_{0}-2\mbox{Re}(\Gamma_{2g,g2,\alpha}^{P_{\alpha}=0})\rho_{2}\right)\right.$
-- ${\cos\theta}\left[2\mbox{Re}(\Gamma_{g0,0g,e0})\rho_{0}+2\mbox{Re}(\Gamma_{2g,e2})\rho_{2}-(\mbox{Re}(\Gamma_{e0,0g})+\mbox{Re}(\Gamma_{e2,2g}))(1-\rho_{0}-\rho_{2})\right],$
$\langle\dot{S}_{y}\rangle$ $=$ $-\sum_{v=0,2}2\mbox{Im}(\Gamma_{ev,vg})(\langle S_{z}\rangle\cos\theta+\langle S_{x}\rangle\sin\theta)/\hbar-2\sum_{v=0,2}\mbox{Re}(\Gamma_{gv,vg}^{P_{\alpha}=0})\langle S_{y}\rangle/\hbar$
-- $\left(\Delta_{ge}-2\sum_{\alpha}P_{\alpha}\cos\theta(\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0})-\mbox{Im}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0})\right)\left(\langle S_{z}\rangle\sin\theta-\langle S_{x}\rangle\cos\theta\right)/\hbar.$
Doing the same for the equations of the empty and doubly state yield:
$\hbar\dot{\rho}_{0}$ $=$ $2\mbox{Re}(\Gamma_{g0,0g}^{P_{\alpha}=0})(1-\rho_{0}-\rho_{2})+4\sum_{\alpha}\mbox{Re}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0})P_{\alpha}\cos\theta(\langle S_{z}\rangle\cos\theta+\langle S_{x}\rangle\sin\theta)/\hbar$
$+$ $4\mbox{Re}(\Gamma_{e0,0g})(\langle S_{z}\rangle\sin\theta-\langle S_{x}\rangle\cos\theta)/\hbar-4\mbox{Re}(\Gamma_{0g,g0}^{P_{\alpha}=0})\rho_{0},$
$\hbar\dot{\rho}_{2}$ $=$ $2\mbox{Re}(\Gamma_{g2,2g}^{P_{\alpha}=0})(1-\rho_{0}-\rho_{2})-4\sum_{\alpha}\mbox{Re}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0})P_{\alpha}\cos\theta(\langle S_{z}\rangle\cos\theta+\langle S_{x}\rangle\sin\theta)/\hbar$
$+$ $4\mbox{Re}(\Gamma_{e2,2g})(\langle S_{z}\rangle\sin\theta-\langle S_{x}\rangle\cos\theta)/\hbar-4\mbox{Re}(\Gamma_{2g,g2}^{P_{\alpha}=0})\rho_{2}.$
By considering that $\sum_{\alpha}P_{\alpha}\left[\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}-\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0})\right]\sin\theta=\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg}(t))$, one sees that the second and third lines in the $S_{x}$ equation, both proportional to the polarization, cancel out, while in the $S_{z}$ equation, we obtain the non-equilibrium spin accumulation. At the same time, this relation further simplifies the equations for the transient state by merging the second and third terms in both equations, ultimately leading to Eqs. LABEL:spin_equations.